### v1

# DitHub: A Modular Framework for Incremental Open-Vocabulary Object Detection

Abstract
Open-Vocabulary object detectors inherently generalize to an unrestricted set of categories, enabling recognition through simple textual prompting. However, adapting these models to rare classes or reinforcing their abilities on specialized domains remains essential. While recent methods rely on monolithic adaptation strategies with a single set of weights, we embrace modular deep learning. We introduce DitHub, a framework designed to build and maintain a library of efficient adaptation modules. Inspired by Version Control Systems, DitHub manages expert modules as branches that can be fetched and merged as needed. This modular approach allows us to conduct an in-depth exploration of the compositional properties of adaptation modules, marking the first such study in Object Detection. Our method achieves state-of-the-art performance on the ODinW-13 benchmark and ODinW-O, a newly introduced benchmark designed to assess class reappearance. For more details, visit our project page: https://aimagelab.github.io/DitHub/
1 Introduction
Object detection is a cornerstone of computer vision, with recent advancements [42, 1, 51, 26, 20] enabling architectures that integrate Vision and Language large-scale pre-training [21]. This has given rise to Open-Vocabulary Object Detection, where models can classify and localize objects beyond a predefined set of categories. The ability to process arbitrary textual queries and detect corresponding objects has profound implications for real-world applications, including medical imaging [15], safety and security [35], and industrial quality control [32]. For instance, the ability to detect rare or atypical objects from a safety camera is crucial for issuing early warnings about uncommon dangers. However, this requires Open-Vocabulary detectors to adapt effectively, enabling them to recognize categories beyond their pre-training knowledge distribution.
Figure 1: DitHub: a novel system to share and update deep adaptation modules for Open-Vocabulary Object Detection. [^2] The creature in the red box is an axolotl, symbolizing rarity as a deviation from pre-training. We use a stylized axolotl as the icon for DitHub.
Recent approaches, such as those detailed in [2], have extended vision-language detectors to incrementally accommodate new categories while preserving robust zero-shot capabilities [34]. However, these methods often employ monolithic adaptation, where all newly acquired knowledge is condensed into a single set of weights. This approach presents challenges in real-world scenarios that demand updates to specific concepts which may reappear under varied input modalities. For instance, in safety and security applications, the class “person” may need to be detected not only in standard RGB imagery but also in thermal imagery, necessitating continuous model adaptation to ensure consistent performance across different data types. Additionally, for rare or complex categories, incremental adaptation is crucial to refine the base model for fine-grained concepts. In such scenarios, a monolithic deep architecture faces challenges similar to those of managing a complex program written on a single page of code. Namely, when treating multiple classes within a unique set of weights, it becomes hard to update specific concepts selectively, without compromising information related to other concepts. Considering less-represented categories, the corresponding concepts can become blurred within the unique set of weights, leading to potential degradation in performance over time.
We advocate for modular approaches, as outlined in Modular Deep Learning [30], which reimagine neural networks as adaptable and composable entities rather than rigid, monolithic structures. By maintaining a library of expert modules, neural architectures can achieve class-specific specialization, activating only the relevant module when needed. Moreover, this modular design enables selective updates and seamless composability, allowing class-specific modules trained in different domains to be efficiently merged. Finally, modularity offers a practical solution to mitigate biases — if a module exhibits undesirable biases toward a particular category, it can be selectively removed or adjusted without disrupting the entire model, ensuring more precise and fair adaptation.
With modularity at its core, we introduce DitHub (where D stands for Detection), a framework inspired by Version Control Systems that efficiently manages a growing library of detection modules. This library expands incrementally as new groups of categories, referred to as tasks [41], are introduced. Upon encountering a new task, DitHub initializes an efficient module (e.g., LoRA [11]), which first undergoes a broad pre-tuning phase across all classes of the task, referred to as the warmup phase. This stage establishes a solid and class-agnostic foundation, enabling the subsequent specialization of the module for specific classes. Following this phase, the module branches into specialized experts, each optimized independently. To enable selective updates for classes previously introduced to the model, DitHub first performs a fetch operation to access potentially existing class-specific modules. If these exist, instead of starting the training from the warmup, DitHub uses a merged initialization — combining the retrieved pre-existing class-specific modules with the current warmup module. This merged initialization facilitates smooth domain knowledge transfer and ensures continuity of class-specific learning across tasks.
Moreover, to address potential memory constraints, DitHub employs a subset of shared learnable parameters across different adaptation modules, enhancing efficiency without compromising individual class adaptability. At inference, it enables the selective activation of specialized modules for rare or complex classes, ensuring robust adaptation while maintaining the backbone’s zero-shot and open-vocabulary capabilities. A high-level overview of our approach is depicted in Footnote 2.
In the context of [2], we evaluate DitHub on the ODinW-13 [17] benchmark, achieving state-of-the-art performance within the Incremental Vision-Language Object Detection paradigm. To further assess our proposal, we introduce ODinW-O (Overlapped), a curated subset of ODinW-35 [16] that focus on classes re-appearing across multiple tasks. Our simple yet effective methodology — without additional loss terms or architectural modifications — allows us to investigate the compositional capabilities of efficient modules, marking the first exploration of this kind in Object Detection. In summary, our key contributions are:
  • We introduce DitHub, a framework inspired by Version Control Systems, designed to manage a growing library of efficient modules for Open-Vocabulary detection.
  • We achieve state-of-the-art performance on ODinW-13 and the newly introduced ODinW-O.
  • We conduct an in-depth analysis of the compositional and specialization capabilities of class-specific modules, a first in the context of Object Detection.
2 Related Works
Incremental Vision-Language Object Detection
Vision-Language Object Detection (VLOD), also referred to as Open-Vocabulary detection [2], is the task of detecting and recognizing objects beyond a fixed set of predefined categories. Early approaches leverage powerful visual and textual encoders (e.g., CLIP [34]) to enable object detectors to recognize virtually any object specified by a textual query [8, 49]. A fundamental trend in this field has been the reliance on large-scale pre-training on extensive datasets [37], which has become standard practice for tasks such as classification. In the context of Vision-Language Object Detection, state-of-the-art methods have followed this paradigm, with GLIP [17] being one of the first notable examples. GLIP reframes object detection as a phrase-grounding problem, leading to a pre-trained model capable of generalizing to unseen objects by integrating textual and visual semantics from large-scale datasets. Building on GLIP’s problem formulation, Grounding DINO [21] extends this paradigm by incorporating merged visual and textual semantics at multiple network stages.
Recently, Incremental Vision-Language Object Detection (IVLOD) [2] has emerged as a natural extension of the VLOD paradigm. By incorporating Incremental Learning [41, 14, 36, 45], IVLOD addresses the challenge of fine-tuning a pre-trained Vision-Language model on specific object categories while preserving its zero-shot [29] capabilities. Following the approach of [2], we adopt Grounding DINO as our backbone and utilize IVLOD as the foundation to explore the merging of class-specific modules.
Parameter Efficient Fine-Tuning and Model Merging
The rise of Transformer models [42, 5] has made pre-training a standard approach in machine learning. Parameter-Efficient Fine-Tuning (PEFT) methods, such as Prefix Tuning [18], Adapters [10], and LoRA [11], adapt pre-trained models to downstream tasks by updating only a small fraction of parameters while keeping the rest frozen.
In the context of Incremental Vision-Language Object Detection, ZiRa [2] introduces a distinct approach that can be viewed as a specialized form of parallel adapters [23], incorporating reparameterizable side branches. In contrast, our work adopts LoRA and explores the potential of merging class-specific LoRA modules.
Several studies have investigated LoRA module fusion. An early approach relies on simple arithmetic operations such as addition and subtraction [50]. More advanced techniques focus on learning optimal combination coefficients, determining the contribution of each LoRA module in the merging process [12, 47, 46, 48]. Recently, researchers have introduced closed-form solutions for LoRA merging [38] and proposed regularization objectives to ensure that the fused modules remain close to the pre-trained weight distribution [31]. In this respect, our approach takes a novel direction, being the first to study module merging in the context of object detection.
3 Preliminaries
Problem setting. We consider a sequence of tasks $\{1,\dots,T\}$, where each task $t$ corresponds to a dataset $D_{t}$. Each dataset consists of samples $(x,y)$, where $x$ represents the input data, and $y$ contains the associated labels, including ground truth bounding boxes.
Given a pre-trained Vision-Language detector $f(\cdot;\Theta)$, parameterized by $\Theta$, the objective of Incremental Vision-Language Object Detection (IVLOD) is to improve performance on newly introduced tasks while preserving original zero-shot capabilities. Formally, for a given task $t$, we seek an adaptation $\Delta$ that minimizes the following objective:
$\min_{\Delta}\mathbb{E}_{(x,y)\sim D_{t}}\;\ell(f(x;(\Theta,\Delta)),\,y)+\ell_{\text{zero-shot}},$ (1)
where $\ell$ is the object detection loss function. The term $(\Theta,\Delta)$ represents the set of adapted model parameters. Additionally, $\ell_{\text{zero-shot}}$ quantifies the performance difference in the zero-shot setting before and after adaptation.
Each task $t$ introduces a set of classes $\mathcal{C}_{t}$. Unlike standard incremental learning [41, 45, 44, 40], IVLOD allows class overlap between tasks. In other words, there exist task pairs $(t,t^{\prime})$, with $t^{\prime}\neq t$, such that $\mathcal{C}_{t}\cap\mathcal{C}_{t^{\prime}}\neq\emptyset$.
LoRA. Low-Rank Adaptation [11] is a technique to significantly reduce the number of trainable parameters. Given a generic layer with weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA introduces low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ — with $r\ll\min(d,k)$ — such that the adapted weight $W^{\prime}$ is:
$W^{\prime}=W+\Delta W,\quad\text{where}\quad\Delta W=BA.$ (2)
The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to fine-tuning the entire model.
4 DitHub
Figure 2: Overview of DitHub. Through three primitives — i.e., branch, fetch, merge — DitHub allows to update a library of class-specific modules for Open-Vocabulary Object Detection, ensuring a balanced integration of both past and present knowledge.
Version Control Systems are ubiquitous tools in computer science, renowned for their capacity to efficiently manage code repositories. Inspired by this capability, we introduce DitHub— a framework that operates akin to a version control system, facilitating the creation and continuous expansion of a library of efficient adaptation modules for vision-language object detectors. To develop and maintain class-specific expert modules, DitHub employs a straightforward yet effective strategy that requires no architectural modifications or additional loss terms, as detailed in Sec. 4.1.
In line with the authors of [2], we consider the pre-trained Grounding DINO as the base architecture for Open-Vocabulary detection. We employ LoRA as a fine-tuning strategy, proposing a novel and efficient pipeline that handles its low-rank matrices with two specific objectives. In our approach, matrix $A$ is dedicated to class-specificity, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency, as further discussed in Sec. 4.2. Finally, when a specialized task arises or more fine-grained adaptation is required (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Sec. 4.3. The pseudocode of DitHub is provided in Alg. 1.
4.1 Training Class-Specific Modular Detectors
To manage a library of class-specific modules, the approach is to train a tailored $A$ matrix for each class, allowing it to specialize accordingly. On the contrary, the matrix $B$ is shared across classes and continually fine-tuned while tasks progress. Thus, at task $t$, we train a total of $|\mathcal{C}_{t}|+1$ matrices for each layer that requires adaptation. The total includes a matrix for each of the $|\mathcal{C}_{t}|$ classes introduced in the current task, along with a single matrix for $B$.
However, in object detection, creating class-specific experts is non-trivial, as a single image may contain multiple bounding boxes spanning different categories, making it difficult to focus on a certain class. To tackle this, we adopt a stochastic training strategy for specializing the $A$ matrices: given an image with multiple objects from different classes, we randomly select one of the present classes and execute a training step on its corresponding class-expert module. However, in our preliminary experiments, this strategy proved ineffective for rare categories, which may appear only a few times in real-world scenarios. Relying solely on random selection prevents the $A$ matrix of rare classes from accessing enough training data.
As better highlighted in Alg. 1, we overcome this issue by introducing an earlier warmup phase at the beginning of each task. This phase is meant to provide a robust, class-agnostic initialization, ensuring that the subsequent specialization stage is more effective.
Warmup. At the onset of the new task, a single and shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Fig. 2, yellow arrow) to capture class-agnostic yet task-specific knowledge. This phase, spanning a predefined number of epochs, establishes a common initialization for the $A$ modules, which are set to undergo specialization in the subsequent stage. By doing so, the specialized modules start from a common base point, a factor that has been shown to greatly enhance model compositionality [27, 33]. Often associated with improved linear mode connectivity between models [7], these properties are leveraged by our approach to select and combine class-specific modules based on input textual queries, as detailed in Sec. 4.3. Moreover, considering rare and underrepresented classes with a few training examples, the warmup phase seeks to establish an already robust and effective initialization.
Specialization. Once the warmup phase is finished, we begin with the specialization phase, the cornerstone of DitHub. To train class-specific modules, DitHub generates $|\mathcal{C}_{t}|$ distinct branches (branch in Fig. 2), each dedicated to a single class. These branches undergo independent optimization, represented by a blue arrow in Fig. 2, to capture class-specific capabilities. In doing so, for each training image, we employ the stochastic strategy outlined at the beginning of Sec. 4.1 to randomly select a single expert for training.
Notably, the modular design of our approach effectively handles potential class overlaps across consecutive tasks. In real-world scenarios, classes may reappear in new domains with altered distributions or semantics. For instance, consider the class “dog”, originally learned from real-world camera images, later appearing in a task involving thermal imagery. The goal is not only to retain prior knowledge but also to develop a “dog” expert capable of recognizing the entity in both domains, accommodating any semantic shifts. In this case, the modularity of DitHub enables the selective adaptation of the “dog” module, while minimizing interference with other classes, such as the “axolotl” module.
To favor selective updates, DitHub performs a fetch operation (Fig. 2) to retrieve any previously stored version of the specialized module, such as $A_{\text{dog}}^{\text{old}}$. If found, this module is merged (merge in Fig. 2) with the current task’s warmup matrix $A_{\text{wu}}$. For instance, considering the class “dog”, the merge operation follows a weighted strategy:
$A_{\text{dog}}^{\text{cur}}=(1-\lambda_{A})A_{\text{dog}}^{\text{old}}+\lambda_{A}A_{\text{wu}}$ (3)
Here, $\lambda_{A}$ is set to a low value, prioritizing consolidated knowledge over new insights from the warmup phase. Details on the impact of $\lambda_{A}$ are in Sec. 5.3. Crucially, our fetching process creates an on-the-fly initialization that blends historical class-specific knowledge with new domain information.
Algorithm 1 DitHub Training
Input: $T$ disjoint tasks each containing $|\mathcal{C}_{t}|$ classes; library $\mathcal{A}\leftarrow\{\}$ of class experts; global $B$.
for task $t\in\{1,\dots,T\}$ do
     $(A_{\text{wu}},B^{\text{opt}})\leftarrow$warmup (training)
     for class $c\in\mathcal{C}_{t}$ do $\triangleright$  Branch
         if $A_{c}\in\mathcal{A}$ then $\triangleright$  Fetch
              $A_{c}^{\text{cur}}\leftarrow$ use Equation 3 $\triangleright$  Merge
         else
              $A_{c}^{\text{cur}}\leftarrow A_{\text{wu}}$
         $(A_{c},B^{\text{opt}})\leftarrow$specialization (training)
         $\mathcal{A}\leftarrow\mathcal{A}\cup A_{c}$
     $B\leftarrow$ use Equation 4
4.2 Bridging Modularity with Memory Efficiency
Figure 3: Memory requirements as the number of classes grows.
To keep our library memory-efficient, we maintain a single shared $B$ matrix instead of a distinct $B$ for each class. As shown in Fig. 3, using class-specific $(A,B)$ pairs results in a linear increase in memory consumption as the number of classes grows. Sharing the $B$ matrix reduces this requirement by half  [^3] In Sec. A.2, we present DitHub’s results when $B$ is not shared.. Indeed, considering DitHub “$B$ shared” (line [Uncaptioned image]), the memory footprint of our approach is consistently lower than the storage needed for the Grounding DINO encoder model alone (line [Uncaptioned image]). This makes DitHub’s memory demands manageable for practical applications. Also, the memory efficiency is further optimized by fine-tuning only a limited number of layers (see App. A).
Due to the shared design, the matrix $B$ is more susceptible to interference and catastrophic forgetting [25]. Consequently, we implement a tailored mechanism that ensures smooth updates as tasks progress. Specifically, at the end of each task $t$, we update the matrix $B$ by merging the previous checkpoint, $B_{t-1}$, with the weights learned after the current task, $B^{\text{opt}}$. This process is formulated as follows:
$B_{t}=(1-\lambda_{B})B_{t-1}+\lambda_{B}B^{\text{opt}}.$ (4)
Unlike Eq. 3, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state. This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, which will be further explained in Sec. 5.3.
4.3 Detection with Class-Specific Modules
With our library of efficient, class-specific modules in place for adapting the Vision-Language backbone, two disjoint inference routes are available for users. Given a textual query specifying the classes of interest, our model tokenizes the prompt and searches token by token to determine if corresponding expert modules are present in the library. If there are correspondences, those relevant modules are fused using a simple average to create a single composite adaptation module. Conversely, for requests targeting individual classes where fine-grained adaptation is desired, the corresponding module can be directly selected and deployed. In our experimental studies, we present results and discussions that address both scenarios.
5 Experimental Studies
This section evaluates DitHub in two parts. The first assesses its adaptability to multiple tasks using the Incremental Vision-Language Object Detection (IVLOD) protocol [2] in Sec. 5.1. The second, in Sec. 5.2, examines the method’s specialization capabilities. The section ends with ablation studies in Sec. 5.3.
5.1 Incremental Vision-Language Object Detection
Shots Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
0 G-Dino 47.41 46.80 19.11 20.82 64.75 59.98 25.34 56.27 54.80 65.94 22.13 62.02 32.85 70.38 57.07
1 TFA 18.84 39.50 18.25 15.81 63.90 50.79 28.47 50.37 29.49 59.16 21.90 50.67 19.86 60.85 43.97
iDETR 44.61 49.82 22.81 23.24 69.75 61.43 31.73 56.27 55.40 62.44 28.45 60.33 43.33 73.64 58.84
AT 44.11 46.23 21.55 23.62 66.60 58.96 27.68 53.97 54.58 62.47 26.94 53.17 20.37 70.31 60.71
ZiRa 46.44 48.56 20.34 19.64 69.47 60.00 30.09 56.27 58.12 65.13 25.83 55.69 39.50 72.59 58.55
DitHub 46.40 49.19 21.33 24.34 69.23 61.99 42.10 57.08 58.85 55.24 24.62 57.58 33.56 70.99 62.50
5 TFA 26.43 45.76 21.92 22.30 67.40 60.72 30.63 53.56 46.80 63.60 26.88 56.26 28.00 64.28 52.49
iDETR 43.51 51.65 25.69 25.53 70.42 62.98 49.98 50.54 54.85 64.80 33.24 57.64 42.36 76.51 56.92
AT 43.67 47.16 14.63 24.97 66.56 64.19 38.85 42.03 55.49 65.20 27.48 52.68 32.21 71.23 57.54
ZiRa 45.41 51.77 24.22 26.40 71.34 63.96 49.54 60.51 60.11 64.39 32.20 56.31 40.61 69.16 54.30
DitHub 45.86 52.85 26.84 28.91 68.46 60.25 51.93 55.10 60.43 59.44 33.18 68.12 38.05 75.04 61.34
10 TFA 34.35 46.61 21.17 22.16 66.82 60.63 32.35 50.15 55.54 64.98 27.59 57.40 28.14 66.82 52.11
iDETR 43.54 53.29 25.39 27.70 65.62 67.58 47.99 60.20 56.32 63.93 35.18 59.37 53.63 74.70 55.12
AT 43.06 47.34 18.73 25.42 69.77 66.34 35.84 48.25 53.39 64.07 28.89 50.33 32.50 66.57 55.37
ZiRa 46.14 53.20 25.56 25.96 71.72 65.28 48.22 62.81 61.24 65.47 35.58 55.73 46.86 70.97 56.14
DitHub 46.54 54.43 26.19 30.28 68.11 65.73 52.82 56.82 63.36 61.85 36.68 66.82 41.46 76.00 61.48
Full TFA 30.97 47.93 23.80 30.65 67.21 61.77 30.52 50.23 47.73 60.91 29.25 61.72 31.42 66.23 61.61
iDETR 37.32 58.71 32.64 46.65 70.99 68.56 55.32 58.88 64.48 71.01 50.33 63.30 39.19 77.12 64.80
AT 42.30 51.14 23.62 39.90 72.32 65.51 31.47 50.48 60.51 66.07 39.09 53.50 34.04 68.07 60.23
OW-DETR 31.22 55.58 28.46 43.78 70.54 67.78 43.84 56.75 63.13 69.51 45.16 58.99 36.99 74.42 63.20
CL-DETR 32.15 57.26 29.35 45.15 71.94 69.90 45.21 58.52 65.10 71.68 46.58 60.83 38.14 76.74 65.18
ZiRA 46.26 57.98 31.76 47.35 71.77 64.74 46.53 62.66 66.39 71.00 48.48 63.03 41.44 76.13 62.44
DitHub 47.01 62.19 34.62 50.65 70.46 68.56 49.28 65.57 69.58 71.10 56.65 70.88 52.82 79.30 68.18
Table 1: Comparison of mAP values across ODinW-13 tasks. Avg represents the average across the $13$ tasks, while zero-shot performance is reported in the ZCOCO column. Best results are highlighted in bold.
Evaluation settings. Following [2], we employ ODinW-13 [17] to assess the model ability to learn new tasks while mitigating forgetting of previously seen data. ODinW-13 consists of $13$ sub-datasets, spanning from the traditional Pascal VOC [6] to more challenging domains with significant distribution shifts, such as thermal imagery. Also, the IVLOD setting mandates that the Open-Vocabulary detector maintains robust zero-shot capabilities after fine-tuning. These capabilities are tested on MS COCO [19].
One issue of ODinW-13 is that most classes are confined to a single task among the $13$. However, DitHub is conceived to handle classes that reoccur across different domains, necessitating updates to their respective modules. To provide an additional perspective on incremental object detection and model reuse, we introduce a variant of ODinW-13, referred to as ODinW-O. Starting from ODinW-35, we construct ODinW-O by retaining only the classes that are shared across tasks while discarding those that appear in only one task. Further details are provided in App. B.
For comparisons, we employ mean Average Precision (mAP) and compare DitHub against six state-of-the-art methods. CL-DETR [22] and OW-DETR [9] are explicitly designed for incremental learning, while TFA [43] specializes in few-shot adaptation. AT [10] has been extended to both incremental and few-shot learning, whereas iDETR [4] integrates both aspects. Finally, ZiRa [2], our primary competitor, is specifically tailored for IVLOD.
Implementation details. We focus on Grounding DINO and adapt only its encoder network, as further validated in Sec. A.1. The merging coefficient $\lambda$ for the $A$ matrix is set to $0.3$ for ODinW-13 and $0.1$ for ODinW-O; for the matrix $B$, we set $\lambda=0.7$. The choice of $\lambda$ and its underlying rationale are further discussed in Sec. 5.3. During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — App. C for further hyperparameter details.
Performance on ODinW-13. Table 1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO). We evaluate performance in both the standard setting with the full dataset and the few-shot scenarios, considering 1, 5, and 10 shots per class. All baseline results, except those of Grounding DINO, ZiRa, and our method, are taken directly from [2]. Each number represents the average of three runs with different random seeds, with standard deviations provided in Sec. D.1.
In the full-shot setting, DitHub outperforms its main competitor, ZiRa, by a substantial + 4.21 mAP points and achieves the best results in nine out of thirteen tasks. Notably, on ZCOCO, our approach outperforms ZiRa by a mAP gain of + 0.75, setting a new state-of-the-art in both incremental and zero-shot retention capabilities. Although DitHub is not explicitly designed for data-scarce conditions, DitHub achieves state-of-the-art performance even in the few-shot setting. In particular, in the five-shot and ten-shot settings, DitHub emerges as the best-performing approach, with the ten-shot scenario showing a slightly larger improvement over ZiRa (+ 1.23 mAP). Other baselines struggle to remain competitive in this regime. Our method also performs well in the one-shot setting, where iDETR achieves the highest average performance, followed closely by DitHub and ZiRa.
These findings — optimal performance in incremental learning, effective retention of zero-shot capabilities, and robustness in the low-data regime — clearly highlight the advantages of our modular approach, with the advanced update system that will be dissected in the analyses below.
ZCOCO Avg Ae Hw Pv Sd Th Ve
G-Dino 47.41 53.15 45.12 67.54 58.11 25.84 70.40 51.87
ZiRa 44.43 57.63 39.92 68.00 64.90 46.26 77.26 49.47
DitHub 46.51 62.38 53.35 71.07 71.01 41.75 80.21 56.90
Table 2: mAP values on ODinW-O. Best in bold.
Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
G-Dino 47.41 46.80 19.11 20.82 64.75 59.98 25.34 56.27 54.80 65.94 22.13 62.02 32.85 70.38 57.07
EnE 46.86 60.96 40.51 44.88 69.67 67.54 51.62 63.42 70.32 70.29 53.76 69.93 45.19 78.36 66.95
DitHub 47.01 62.19 34.62 50.65 70.46 68.56 49.28 65.57 69.58 71.10 56.65 70.88 52.82 79.30 68.18
Table 3: Experts not Experts (EneE). Training by selecting a random expert.
Performance on ODinW-O. The ODinW-O scenario emphasizes that frequently encountered classes may reappear in subsequent tasks or at later times, such as the class “person” in common industrial and smart-city scenarios. As detailed in Sec. B.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks. In this benchmark, we primarily compare our approach with ZiRa, as it is the most competitive method in the standard IVLOD setting, excluding ours. The results in Tab. 2 indicate that our approach secures a substantial improvement of + 4.75 in mAP and a + 2.08 mAP gain on ZCOCO over ZiRa. We attribute the success of our method to its class-oriented modular design, which enables selective updates to recurring concepts. This design effectively prevents the model from inadvertently overwriting crucial knowledge associated with other classes, tasks, or domains.
5.2 Evaluation of Specialized Detection Modules
Here, we delve into the properties of our specialization framework and the benefits of composing class-specific experts. To this aim, we introduce an ablative approach called Experts-not-Experts (EnE), which emphasizes the absence of class specialization in our experts. EnE is identical to DitHub, except during training, where it does not employ the stochastic strategy introduced in Sec. 4.1 and designed to specialize the experts. Instead, given an image from the current task, the EnE approach performs the training step on one module chosen at random among the classes of the entire task. As a result, modules no longer specialize in a class but instead learn from the whole task data, including images without their designated class. Consequently, they become general task learners rather than class-specific experts.
Figure 4: Specialization capabilities for rare classes of ODinW-13.
In Tab. 3, we offer a comparative analysis between our DitHub and its non-specialized variant, i.e., EnE. Notice that, at inference time, EnE acts as an ensemble of modules, each trained on the same data but with a different random ordering. In contrast, DitHub employs class-specialized modules with an expertise on data containing the same class. This targeted specialization improves performance, as shown in Tab. 3, where per-class specialization results in a + 1.23 mAP improvement on average. Additionally, albeit to a lesser extent, the ZCOCO metric still benefits from an improvement. These results affirm that a system composed of modular experts, each a custodian of distinct knowledge, offers a promising approach for systems requiring continuous and selective adaptation.
Evaluation on rare classes. As discussed in Section 4.3, DitHub facilitates fine-grained control during inference, enabling users to tailor the system to their specific use case and deploy the most appropriate modules. This includes the option to use a single module specialized on a target class. To assess single-module effectiveness, we conduct a targeted evaluation on rare classes from ODinW-13, activating only the corresponding target class module. Results in Fig. 4 show that DitHub consistently achieves the best results across rare classes, highlighting the benefits of class specialization. EnE and DitHub outperform the zero-shot Grounding DINO baseline, though EnE underperforms for the Starfish category.
Figure 5: Showing the natural unlearning capabilities of DitHub.
Specialized modules allows training-free unlearning. The advantages of DitHub extend beyond additive incremental knowledge compositionality. Indeed, we can show that expert modules can be useful to undermine or remove some specific knowledge or content from a model, by simply subtracting the corresponding weights from the base model. This operation, which resembles the concept of zero-training unlearning, has been explored by the authors of [13, 28] in the context of image and text categorization tasks; to the best of our knowledge, we are the first to explore its application within the context of object detection.
Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class module from the pre-trained model: $\mathbf{W}^{\prime}=\mathbf{W}-\alpha BA_{c}$, where we set $\alpha=0.3$. As a result, we provide a qualitative visualization shown in Fig. 5: on the left, the zero-shot model correctly recognizes common classes, while on the right, applying the negative module of a specific class reduces its mAP, effectively unlearning it and leading to false negatives. Additionally, we provide a quantitative evaluation of the unlearning process in Sec. D.2.
5.3 Ablation Studies
Figure 6: Sensitivity analysis on $\lambda_{A}$ and $\lambda_{B}$. Performance measured as harmonic mean of Avg mAP (across tasks) and ZCOCO.
Sensitivity Analysis of $\lambda.$ Analyzing the variation of the merging coefficients $\lambda_{A}$ and $\lambda_{B}$ offers valuable insights into our methodology. To show that, we define an aggregate metric that is the harmonic mean between DitHub’s full-shot performance on ZCOCO and Avg, ensuring a balanced evaluation of specialization (Avg) and zero-shot generalization (ZCOCO). Figure 6 presents the performance trends for ODinW-13 (top) and ODinW-O (bottom) while varying $\lambda_{A}$ and $\lambda_{B}$ in increments of $0.2$.
Considering $\lambda_{A}$, its role is significant while merging the expert A matrices (see Eq. 3). From Fig. 6, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective. Interestingly, while $\lambda_{A}=0.3$ performs best for ODinW-13, ODinW-O exhibits a monotonic decline as $\lambda_{A}$ increases. This contrast arises from dataset structures: ODinW-O consists solely of overlapping classes across tasks, where retaining past expertise is critical, whereas ODinW-13 features diverse class distributions per task, requiring greater integration of domain-specific warmup knowledge.
For the incremental adaptation of the shared $B$ matrix, a different trend emerges. The harmonic mean increases with $\lambda_{B}$, peaking at $0.7$ for both datasets. This suggests that prioritizing the knowledge in the $B^{\text{opt}}$ matrix from the current task, as shown in Eq. 4, is crucial. Our intuition, supported by [24], is that the latest $B^{\text{opt}}$ matrix integrates knowledge from both earlier and recent tasks, making it more valuable than $B_{t-1}$, which is trained solely on earlier tasks.
Figure 7: Ablative study on the DitHub components. Results are expressed as average performance across the tasks of ODinW-13.
On Warmup and Shared B. To evaluate the impact of each component in DitHub, we progressively integrate them and present the results in Fig. 7. We start with the zero-shot performance of Grounding DINO (GD) as a reference, followed by a baseline fine-tuning procedure (Base in Fig. 7), where an $A$ matrix is trained for each class without merging strategies, omitting the warmup phase and using a shared $B$ matrix. This method leads to complete forgetting as new tasks are introduced, but still outperforms Grounding DINO, which struggles with pretraining knowledge alone.
Building on this baseline, we first incorporate Eq. 4, but see no significant performance gain. This is expected, as the shared $B$ matrix only represents a tiny portion of the total parameters. Without class-specific focus, Eq. 4 alone does not enhance retention. However, adding the warmup phase (+wu) provides a clear improvement, as it stabilizes training and helps rare classes by offering a robust initialization.
When we further include Eq. 4 in the warmup-enhanced fine-tuning strategy, a modest improvement is observed. However, the most substantial performance gain occurs when Eq. 3 is also integrated into the methodology (DitHub in Fig. 7). This highlights the importance of preserving past domain knowledge within class-specific modules, ensuring a more stable and effective incremental learning process. Notably, the combination of all three key components — the warmup phase for reliable initialization, the merging strategy for $A$ matrices to retain class-specific knowledge across domains, and the shared $B$ matrix approach for efficiency — results in a significant + 6.22 improvement over the baseline where none of these strategies are applied.
6 Conclusions
In this research, we introduced DitHub, a methodology akin to a version control system that maintains an expanding library of modules to adapt open-vocabulary object detectors. Our approach achieves state-of-the-art performance in the Incremental Vision-Language Object Detection scenario. The incremental nature of this domain allowed us to explore the compositional capabilities of efficient modules, marking the first such investigation in Object Detection.
Looking ahead, we aim to further leverage DitHub’s modular library management to deepen our understanding of knowledge transfer between efficient modules for object detection. For instance, if our library contains an “axolotl” module trained on real-world images and we possess domain knowledge for the cartoon image domain, can we synthesize a module specialized in detecting cartoon “axolotls” through knowledge transfer? DitHub will provide a fork of this manuscript to explore such questions.
Supplementary Material Overview
This supplementary material complements the main manuscript by providing additional analyses, insights, and results on DitHub. Specifically, we delve deeper into efficiency considerations of our adaptation approach (App. A), highlighting blockwise adaptation and LoRA specializations. We introduce the novel ODinW-O benchmark for incremental learning with overlapping classes and domain shifts (App. B), and present insights on unlearning class-specific knowledge (Sec. D.2). Additionally, we clarify critical implementation specifics (App. C), present extensive robustness checks (Sec. D.1), and showcase qualitative examples (App. E) to illustrate the practical benefits of our methodology.
Appendix A Further Insights on Efficiency
A.1 Blockwise Adaptation Analysis
While Grounding DINO consists of four main components — an image backbone for visual encoding, a textual backbone for language understanding, and an encoder-decoder pair for detection — we focus our adaptation on the latter two. This choice is motivated by the need for a more detection-centric adaptation and efficiency considerations, as the textual backbone, being a BERT Base model [3], accounts for more than half of Grounding DINO ’s parameters. In Table B, we present results for adapting only the encoder (Figure A), only the decoder (Figure A), and both encoder and decoder (Figure A). DitHub employs the first approach — adapting only the encoder — as it delivers the best performance on both ZCOCO and Avg metrics.
A.2 Employing Full LoRA for Class-Specificity
While our methodology leverages a shared and unique $B$ matrix for all tasks, Tab. A presents results obtained using class-specific pairs. In this variant of DitHub, both the $A$ and $B$ low-rank matrices of LoRA undergo the same specialized optimization process: an initial warmup phase followed by dedicated specialization. In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Section 4.1. Applying our full set of strategies to both matrices yields a + 1.03 improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Fig. 3). Prioritizing efficiency, we opt for a lightweight implementation that trades a marginal performance gain for significantly reduced memory requirements. Notably, ZCOCO results favor our efficient approach, as a single, unspecialized $B$ matrix remains more aligned with the pretraining distribution, thereby mitigating degradation in zero-shot capabilities.
Method ZCOCO Avg
DitHub 47.01 62.19
DitHub (LoRA full) 46.66 63.22
Table A: LoRA full results on ODinW-13, with DitHub results taken from Table 1 for reference.
(a) Encoder
(b) Decoder
(c) Decoder/Encoder
Figure A: Blockwise analysis of adaptation strategies: (a) Adaptation of the encoder only (our chosen approach), (b) Adaptation of the decoder only and (c) Joint adaptation of both the decoder and encoder.
Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
DitHub 47.01 62.19 34.62 50.65 70.46 68.56 49.28 65.57 69.58 71.10 56.65 70.88 52.82 79.30 68.18
Only Dec 46.11 54.35 28.00 35.22 65.28 65.11 37.62 59.71 68.11 66.27 35.22 67.34 39.46 73.87 65.30
Dec/Enc 46.55 59.92 32.84 43.16 71.47 67.44 49.78 62.84 70.32 70.83 49.24 70.89 46.12 77.05 67.48
Table B: Performance comparison across datasets when adapting only the decoder (”Only Dec”), both encoder and decoder (”Dec/Enc”), and encoder only (DitHub).
Appendix B Additional Benchmark Details
This section provides additional details regarding the datasets utilized in our evaluation framework. We first outline the key characteristics of ODinW-13 and MS COCO, as these datasets constitute the foundation for evaluating the specialization and zero-shot capabilities of DitHub. Subsequently, we introduce and discuss the ODinW-O benchmark in greater depth, which we explicitly design to evaluate incremental learning performance under scenarios involving recurring categories across multiple tasks, accompanied by domain shifts. By leveraging these datasets, we establish a rigorous evaluation protocol to comprehensively assess model robustness in challenging real-world scenarios.
B.1 ODinW-13 and MS COCO
Following [2], we assess our model specialization using the Object Detection in the Wild (ODinW-13) [17] benchmark and the MS COCO [19] dataset to verify whether the zero-shot capabilities are preserved after the sequential fine-tuning on ODinW-13.
Unlike structured and standardized datasets such as MS COCO, ODinW-13 aggregates data from thirteen sources: Aerial Maritime Drone (Ae), Aquarium (Aq), Cottontail Rabbit (Co), EgoHands (Eg), Mushrooms (Mu), Packages (Pa), Pascal VOC (Pv), Pistols (Pi), Pothole (Po), Raccoon (Ra), Shellfish (Sh), Thermal Dogs and People (Th), Vehicles (Ve). This comprehensive collection covers rare object categories, diverse contextual conditions, and challenging domains. Consequently, the ODinW-13 benchmark offers an ideal scenario to evaluate the capability of DitHub to sequentially adapt to heterogeneous task distributions while effectively mitigating catastrophic forgetting on previously learned tasks.
Following the convention established by [2], we use the term ZCOCO to denote the zero-shot evaluation conducted on the MS COCO validation set comprising $5000$ samples and consisting of annotations for 80 common object categories. The ZCOCO evaluation specifically measures the robustness and generalization capabilities of DitHub in zero-shot scenarios, after training on each successive task from ODinW-13.
B.2 ODinW-O
We introduce a new benchmark, ODinW-O (Overlapped), designed to further evaluate the class-incremental and specialization capabilities of DitHub. Derived from ODinW-35, ODinW-O retains only classes shared across at least one task (overlapping classes), while removing tasks with no shared classes. This filtering is necessary because classes that do not reappear in subsequent tasks present an ideal case in the DitHub framework. Specifically, their corresponding non-recurring class modules remain frozen, requiring no further adaptation. In contrast, ODinW-O eliminates this scenario by ensuring that all retained classes appear in multiple tasks. By enforcing strictly overlapping classes, ODinW-O provides a challenging benchmark for evaluating incremental learning, specifically tailored to assess the robustness of our method. Additionally, to rigorously evaluate the model robustness to domain shifts, we designed ODinW-O by explicitly selecting datasets that feature markedly different contexts (Fig. B), enabling clear insights into how effectively the model adapts to variations in visual appearance, environmental conditions, and imaging domains. We select the following datasets, already presented in Sec. B.1, and extracted only the overlapping classes:
  • Aerial Maritime Drone: boat and car,
  • Pascal VOC: boat, car, dog, and person
  • Thermal Dogs and People: dog and person,
  • Vehicles: car and truck,
Additionally, we introduce two further datasets from ODinW-35 to emphasize domain variation:
  • Hard Hat Workers: person
  • Self Driving Cars: car and truck
Despite ODinW-O enforcing class overlap and introducing domain shifts, our model consistently achieves state-of-the-art performance by effectively refining class knowledge while adapting across diverse contexts.
Person Car
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure B: Visualization of domain shifts for shared categories across different tasks within the ODinW-O benchmark.
Appendix C Implementation Details
All experiments for DitHub were conducted on a single RTX A5000 GPU with 24GB of VRAM, requiring approximately 8 hours per training run with a batch size of 2. During training, our implementation in PyTorch efficiently utilizes around 10 GB of VRAM. We employed GroundingDINO-T (Swin-T) as an open-vocabulary object detector, pretrained on O365 [39], GoldG [17], and Cap4M [17] datasets. We performed training with random shuffling across three different seeds to ensure robustness.
For LoRA fine-tuning, we perform low-rank updates to the attention layers with a rank of 16, applying a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-2}$. This approach enables efficient adaptation while significantly reducing computational overhead and memory storage for each class-specialized module. Optimization is performed using AdamW. We rely solely on the contrastive classification loss and the localization loss of Grounding DINO, without introducing any additional loss functions. The full Grounding DINO model, after integrating 44 classes during training on ODinW-13, consists of approximately 173 million parameters, with 19.6 million LoRA parameters actively updated during training. After fine-tuning on ODinW-13, only 75.65 MB of LoRA parameters are retained for storage.
Appendix D Additional Results
Method ZCOCO Avg Ae Hw Pv Sd Th Ve
ZiRa ±$\pm 0.25$$\pm 0.27$$\pm 1.76$$\pm 1.14$$\pm 1.35$$\pm 0.32$$\pm 1.75$$\pm 1.76$
Ours ±$\pm 0.24$$\pm 0.66$$\pm 2.39$$\pm 1.88$$\pm 0.76$$\pm 1.31$$\pm 0.46$$\pm 0.84$
Table C: Standard deviations of the reproduced methods in Table 2.
D.1 Standard Deviations
Shots Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
1 ZiRa $\pm$ 0.90 $\pm$ 2.05 $\pm$ 1.49 $\pm$ 0.17 $\pm$ 2.16 $\pm$ 0.98 $\pm$ 0.79 $\pm$ 1.63 $\pm$ 1.33 $\pm$ 7.91 $\pm$ 3.29 $\pm$ 12.17 $\pm$ 4.45 $\pm$ 3.07 $\pm$ 3.25
DitHub $\pm$ 0.60 $\pm$ 0.81 $\pm$ 0.95 $\pm$ 0.57 $\pm$ 2.66 $\pm$ 1.58 $\pm$ 3.72 $\pm$ 0.00 $\pm$ 1.53 $\pm$ 1.10 $\pm$ 2.90 $\pm$ 7.03 $\pm$ 3.68 $\pm$ 4.66 $\pm$ 2.46
5 ZiRa $\pm$ 0.23 $\pm$ 0.43 $\pm$ 3.76 $\pm$ 2.15 $\pm$ 1.40 $\pm$ 1.08 $\pm$ 3.54 $\pm$ 3.74 $\pm$ 1.54 $\pm$ 4.26 $\pm$ 3.73 $\pm$ 4.13 $\pm$ 2.03 $\pm$ 2.02 $\pm$ 4.16
DitHub $\pm$ 0.42 $\pm$ 1.27 $\pm$ 2.52 $\pm$ 0.70 $\pm$ 2.42 $\pm$ 6.74 $\pm$ 0.76 $\pm$ 3.27 $\pm$ 0.17 $\pm$ 5.49 $\pm$ 4.79 $\pm$ 2.22 $\pm$ 4.29 $\pm$ 1.81 $\pm$ 1.54
10 ZiRa $\pm$ 0.28 $\pm$ 0.66 $\pm$ 0.44 $\pm$ 2.64 $\pm$ 2.77 $\pm$ 0.52 $\pm$ 6.43 $\pm$ 0.97 $\pm$ 1.04 $\pm$ 2.17 $\pm$ 2.04 $\pm$ 3.90 $\pm$ 1.21 $\pm$ 3.71 $\pm$ 0.87
DitHub $\pm$ 0.20 $\pm$ 0.54 $\pm$ 2.26 $\pm$ 0.72 $\pm$ 4.23 $\pm$ 3.53 $\pm$ 1.35 $\pm$ 3.19 $\pm$ 0.30 $\pm$ 2.74 $\pm$ 0.34 $\pm$ 2.99 $\pm$ 1.01 $\pm$ 1.15 $\pm$ 1.27
Full ZiRA $\pm$ 0.44 $\pm$ 0.68 $\pm$ 2.73 $\pm$ 1.17 $\pm$ 1.39 $\pm$ 4.77 $\pm$ 7.26 $\pm$ 2.19 $\pm$ 1.33 $\pm$ 0.33 $\pm$ 4.14 $\pm$ 3.57 $\pm$ 2.44 $\pm$ 2.82 $\pm$ 3.13
DitHub $\pm$ 0.16 $\pm$ 0.44 $\pm$ 2.39 $\pm$ 1.00 $\pm$ 2.83 $\pm$ 0.48 $\pm$ 1.34 $\pm$ 0.60 $\pm$ 0.38 $\pm$ 1.68 $\pm$ 0.76 $\pm$ 1.57 $\pm$ 2.87 $\pm$ 0.69 $\pm$ 1.01
Table D: Standard deviations of the reproduced methods in Table 1.
We report the standard deviations of the reproduced methods from Tab. 1 and Tab. 2 in Tab. D and C, respectively, computed over three runs with different random seeds.
D.2 Unlearning Additionals
Figure C: Matrix visualization of mAP degradation per class after removing class modules. Colors are normalized column-wise for readability; darker cells indicate greater performance drops.
As discussed in Sec. 5.2, our modular approach enables the precise removal of specific knowledge from Grounding DINO by subtracting corresponding experts. Figure C illustrates this process: each column represents a target class’s mAP after unlearning different modules, with colors normalized column-wise to highlight relative drops. The darker diagonal pattern indicates that removing a class module leads to the most significant performance drop for its corresponding class. This confirms that our modules effectively encapsulate specialized knowledge, as unlearning them selectively impacts their respective classes.
Appendix E Qualitatives
In Fig. D, we compare zero-shot Grounding DINO (left) with ZiRA (center) and DitHub (right). Our method excels at detecting challenging objects, including small ones that other models miss. Additionally, DitHub avoids a false positive (ZiRA) and a false negative (Grounding DINO), demonstrating more reliable predictions. It also correctly identifies both cars (last row) despite occlusions and distractors. This improved stability stems from per-class specialization, which mitigates hallucinations and ensures smoother adaptation to domain shifts.
Grounding DINO ZiRa DitHub
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure D: Qualitative comparison of Grounding DINO (left), ZiRA (center), and DitHub (right).

### v2

@@
 Abstract
-Open-Vocabulary object detectors inherently generalize to an unrestricted set of categories, enabling recognition through simple textual prompting.
-However, adapting these models to rare classes or reinforcing their abilities on specialized domains remains essential.
+Open-Vocabulary object detectors can generalize to an unrestricted set of categories through simple textual prompting.
+However, adapting these models to rare classes or reinforcing their abilities on multiple specialized domains remains essential.
 While recent methods rely on monolithic adaptation strategies with a single set of weights, we embrace modular deep learning.
@@
 Our method achieves state-of-the-art performance on the ODinW-13 benchmark and ODinW-O, a newly introduced benchmark designed to assess class reappearance.
-For more details, visit our project page: https://aimagelab.github.io/DitHub/
 1 Introduction
@@
 However, this requires Open-Vocabulary detectors to adapt effectively, enabling them to recognize categories beyond their pre-training knowledge distribution.
-Figure 1: DitHub: a novel system to share and update deep adaptation modules for Open-Vocabulary Object Detection.
-[^2] The creature in the red box is an axolotl, symbolizing rarity as a deviation from pre-training.
-We use a stylized axolotl as the icon for DitHub.
-Recent approaches, such as those detailed in [2], have extended vision-language detectors to incrementally accommodate new categories while preserving robust zero-shot capabilities [34].
-However, these methods often employ monolithic adaptation, where all newly acquired knowledge is condensed into a single set of weights.
+Recent approaches, such as those detailed in [4], have extended Vision-Language detectors to incrementally accommodate new categories while preserving robust capabilities [39].
+However, these methods employ monolithic adaptation, where all newly acquired knowledge is condensed into a single set of weights.
 This approach presents challenges in real-world scenarios that demand updates to specific concepts which may reappear under varied input modalities.
@@
 Additionally, for rare or complex categories, incremental adaptation is crucial to refine the base model for fine-grained concepts.
-In such scenarios, a monolithic deep architecture faces challenges similar to those of managing a complex program written on a single page of code.
+In such scenarios, a monolithic deep architecture faces challenges similar to managing a complex program written on a single page of code.
 Namely, when treating multiple classes within a unique set of weights, it becomes hard to update specific concepts selectively, without compromising information related to other concepts.
 Considering less-represented categories, the corresponding concepts can become blurred within the unique set of weights, leading to potential degradation in performance over time.
+Figure 1: DitHub: a novel system to share and update efficient adaptation modules for Open-Vocabulary Object Detection.
+[^2] The creature in the red box is an axolotl, symbolizing rarity as a deviation from pre-training.
+We use a stylized axolotl as the icon for DitHub.
 We advocate for modular approaches, as outlined in Modular Deep Learning [35], which reimagine neural networks as adaptable and composable entities rather than rigid, monolithic structures.
@@
 This stage establishes a solid and class-agnostic foundation, enabling the subsequent specialization of the module for specific classes.
-Following this phase, the module branches into specialized experts, each optimized independently.
+Following the warmup phase, the module branches into specialized experts, each optimized independently.
 To enable selective updates for classes previously introduced to the model, DitHub first performs a fetch operation to access potentially existing class-specific modules.
@@
 This merged initialization facilitates smooth domain knowledge transfer and ensures continuity of class-specific learning across tasks.
-Moreover, to address potential memory constraints, DitHub employs a subset of shared learnable parameters across different adaptation modules, enhancing efficiency without compromising individual class adaptability.
+Moreover, to ensure scalability with respect to the number of classes, DitHub shares a subset of parameters across different adaptation modules, enhancing memory efficiency without compromising individual class adaptability.
 At inference, it enables the selective activation of specialized modules for rare or complex classes, ensuring robust adaptation while maintaining the backbone’s zero-shot and Open-Vocabulary capabilities.
 A high-level overview of our approach is depicted in Footnote 2.
-In the context of [2], we evaluate DitHub on the ODinW-13 [17] benchmark, achieving state-of-the-art performance within the Incremental Vision-Language Object Detection paradigm.
-To further assess our proposal, we introduce ODinW-O (Overlapped), a curated subset of ODinW-35 [16] that focus on classes re-appearing across multiple tasks.
-Our simple yet effective methodology — without additional loss terms or architectural modifications — allows us to investigate the compositional capabilities of efficient modules, marking the first exploration of this kind in Object Detection.
+Following [4], we evaluate DitHub on the ODinW-13 [20] benchmark, where it outperforms the state-of-the-art by a substantial $\boldsymbol{\texttt{+}4.21}$ mAP in the Incremental Vision-Language Object Detection setting.
+We also introduce ODinW-O (Overlapped), a curated subset of ODinW-35 [19] focusing on classes re-appearing across multiple tasks.
+Furthermore, we thoroughly investigate and ablate the compositional capabilities of our modular approach, marking the first such exploration in Object Detection.
 In summary, our key contributions are:
+  • We investigate the importance of decoupling a warmup phase from a specialization one in the context of efficient adaptation modules.
   • We introduce DitHub, a framework inspired by Version Control Systems, designed to manage a growing library of efficient modules for Open-Vocabulary detection.
@@
 2 Related Works
-Incremental Vision-Language Object Detection
+Incremental Vision-Language Object Detection.
 Vision-Language Object Detection (VLOD), also referred to as Open-Vocabulary detection [4], is the task of detecting and recognizing objects beyond a fixed set of predefined categories.
@@
 By incorporating Incremental Learning [46, 17, 41, 50], IVLOD addresses the challenge of fine-tuning a pre-trained Vision-Language model on specific object categories while preserving its zero-shot [34] capabilities.
-Following the approach of [2], we adopt Grounding DINO as our backbone and utilize IVLOD as the foundation to explore the merging of class-specific modules.
-Parameter Efficient Fine-Tuning and Model Merging
+Following [4], we adopt Grounding DINO as the backbone and use IVLOD to explore the merging of class-specific modules.
+Parameter Efficient Fine-Tuning and Model Merging.
 The rise of Transformer models [47, 7] has made pre-training a standard approach in machine learning.
 Parameter-Efficient Fine-Tuning (PEFT) methods, such as Prefix Tuning [21], Adapters [12], and LoRA [13], adapt pre-trained models to downstream tasks by updating only a small fraction of parameters while keeping the rest frozen.
-In the context of Incremental Vision-Language Object Detection, ZiRa [2] introduces a distinct approach that can be viewed as a specialized form of parallel adapters [23], incorporating reparameterizable side branches.
+In the context of IVLOD, ZiRa [4] introduces a distinct approach that can be viewed as a specialized form of parallel adapters [27], incorporating reparameterizable side branches.
 In contrast, our work adopts LoRA and explores the potential of merging class-specific LoRA modules.
@@
 An early approach relies on simple arithmetic operations such as addition and subtraction [55].
-More advanced techniques focus on learning optimal combination coefficients, determining the contribution of each LoRA module in the merging process [12, 47, 46, 48].
+More advanced techniques focus on learning optimal combination coefficients, determining the contribution of each LoRA module in the merging process [14, 52, 51, 28, 53].
 Recently, researchers have introduced closed-form solutions for LoRA merging [43] and proposed regularization objectives to ensure that the fused modules remain close to the pre-trained weight distribution [36].
@@
 Problem setting.
-We consider a sequence of tasks $\{1,\dots,T\}$, where each task $t$ corresponds to a dataset $D_{t}$.
+We consider a sequence of tasks $\{1,\dots,T\}$, where each task $t$ corresponds to a dataset $\mathcal{D}_{t}$.
 Each dataset consists of samples $(x,y)$, where $x$ represents the input data, and $y$ contains the associated labels, including ground truth bounding boxes.
@@
 Formally, for a given task $t$, we seek an adaptation $\Delta$ that minimizes the following objective:
-$\min_{\Delta}\mathbb{E}_{(x,y)\sim D_{t}}\;\ell(f(x;(\Theta,\Delta)),\,y)+\ell_{\text{zero-shot}},$ (1)
+$\min_{\Delta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\;\ell(f(x;(\Theta,\Delta)),\,y)+\ell_{\text{zero-shot}},$ (1)
 where $\ell$ is the Object Detection loss function.
@@
 LoRA.
-Low-Rank Adaptation [11] is a technique to significantly reduce the number of trainable parameters.
-Given a generic layer with weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA introduces low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ — with $r\ll\min(d,k)$ — such that the adapted weight $W^{\prime}$ is:
-$W^{\prime}=W+\Delta W,\quad\text{where}\quad\Delta W=BA.$ (2)
-The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to fine-tuning the entire model.
+Given a layer with weight $W\in\mathbb{R}^{d\times k}$, Low-Rank Adaptation [13] (LoRA) uses low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ – with $r\ll\min(d,k)$ – such that the adapted weight $W^{\prime}=W+\Delta W=W+BA$.
+The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to full fine-tuning.
 4 DitHub
-Figure 2: Overview of DitHub.
-Through three primitives — i.e., branch, fetch, merge — DitHub allows to update a library of class-specific modules for Open-Vocabulary Object Detection, ensuring a balanced integration of both past and present knowledge.
-Version Control Systems are ubiquitous tools in computer science, renowned for their capacity to efficiently manage code repositories.
-Inspired by this capability, we introduce DitHub— a framework that operates akin to a version control system, facilitating the creation and continuous expansion of a library of efficient adaptation modules for vision-language object detectors.
-To develop and maintain class-specific expert modules, DitHub employs a straightforward yet effective strategy that requires no architectural modifications or additional loss terms, as detailed in Sec.
-4.1.
-In line with the authors of [2], we consider the pre-trained Grounding DINO as the base architecture for Open-Vocabulary detection.
-We employ LoRA as a fine-tuning strategy, proposing a novel and efficient pipeline that handles its low-rank matrices with two specific objectives.
-In our approach, matrix $A$ is dedicated to class-specificity, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency, as further discussed in Sec.
-4.2.
-Finally, when a specialized task arises or more fine-grained adaptation is required (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Sec.
+Figure 2: Overview of DitHub. Through three primitives – i.e., branch, fetch, merge – DitHub allows to update a library of class-specific modules for Open-Vocabulary Object Detection.
+Version Control Systems are ubiquitous tools used to manage and track changes in code repositories.
+Inspired by this, we propose DitHub — a framework that operates akin to a Version Control System, enabling the creation and expansion of a library of efficient adaptation modules for Vision-Language object detectors.
+DitHub develops and maintains class-specific expert modules using a simple yet effective strategy that requires no architectural modifications or additional loss terms (Sec.
+4.1).
+Following [4], we consider the pre-trained Grounding DINO as the base architecture for Open-Vocabulary detection and use LoRA for fine-tuning.
+We propose an efficient pipeline that handles LoRA’s low-rank matrices with two objectives: matrix $A$ encodes class-specific knowledge, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency (Sec.
+4.2).
+Finally, when fine-grained adaptation is needed (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Sec.
 4.3.
@@
 4.1 Training Class-Specific Modular Detectors
-To manage a library of class-specific modules, the approach is to train a tailored $A$ matrix for each class, allowing it to specialize accordingly.
-On the contrary, the matrix $B$ is shared across classes and continually fine-tuned while tasks progress.
-Thus, at task $t$, we train a total of $|\mathcal{C}_{t}|+1$ matrices for each layer that requires adaptation.
-The total includes a matrix for each of the $|\mathcal{C}_{t}|$ classes introduced in the current task, along with a single matrix for $B$.
-However, in object detection, creating class-specific experts is non-trivial, as a single image may contain multiple bounding boxes spanning different categories, making it difficult to focus on a certain class.
-To tackle this, we adopt a stochastic training strategy for specializing the $A$ matrices: given an image with multiple objects from different classes, we randomly select one of the present classes and execute a training step on its corresponding class-expert module.
-However, in our preliminary experiments, this strategy proved ineffective for rare categories, which may appear only a few times in real-world scenarios.
-Relying solely on random selection prevents the $A$ matrix of rare classes from accessing enough training data.
+To manage a library of class-specific modules, our approach is to train a tailored $A$ matrix for each class to enable specialization.
+In contrast, the $B$ matrix is shared across classes and continuously fine-tuned as tasks progress.
+Thus, at task $t$, we train a total of $|\mathcal{C}_{t}|+1$ matrices per adaptable layer — one for each class and one shared $B$ matrix.
+However, in Object Detection, creating class-specific experts is non-trivial, as images often contain multiple objects from different categories, making it difficult to isolate a single class.
+To tackle this, we adopt a stochastic training strategy for specializing the $A$ matrices: for each image, we randomly select one of the present classes and update its corresponding $A$ matrix.
+However, preliminary experiments show this approach is ineffective for rare categories, as random selection limits their exposure to sufficient training data.
 As better highlighted in Alg.
@@
 Warmup.
-At the onset of the new task, a single and shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Fig.
+At the onset of the new task, a shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Fig.
 2, yellow arrow) to capture class-agnostic yet task-specific knowledge.
-This phase, spanning a predefined number of epochs, establishes a common initialization for the $A$ modules, which are set to undergo specialization in the subsequent stage.
+This phase, run for a fixed number of epochs, provides a common initialization for the $A$ modules before specialization.
 By doing so, the specialized modules start from a common base point, a factor that has been shown to greatly enhance model compositionality [32, 38].
-Often associated with improved linear mode connectivity between models [7], these properties are leveraged by our approach to select and combine class-specific modules based on input textual queries, as detailed in Sec.
+Often associated with improved linear mode connectivity between models [9], these properties are leveraged by our approach to combine class-specific modules based on input textual queries, as detailed in Sec.
 4.3.
@@
 Specialization.
-Once the warmup phase is finished, we begin with the specialization phase, the cornerstone of DitHub.
+After the warmup phase, we proceed to the core of DitHub: the specialization phase.
 To train class-specific modules, DitHub generates $|\mathcal{C}_{t}|$ distinct branches (branch in Fig.
@@
 2, to capture class-specific capabilities.
-In doing so, for each training image, we employ the stochastic strategy outlined at the beginning of Sec.
+In doing so, for each training image, we employ the stochastic strategy introduced in Sec.
 4.1 to randomly select a single expert for training.
@@
 For instance, consider the class “dog”, originally learned from real-world camera images, later appearing in a task involving thermal imagery.
-The goal is not only to retain prior knowledge but also to develop a “dog” expert capable of recognizing the entity in both domains, accommodating any semantic shifts.
+The goal is to retain prior knowledge and enable a “dog” expert to recognize the entity across domains, despite semantic variation.
 In this case, the modularity of DitHub enables the selective adaptation of the “dog” module, while minimizing interference with other classes, such as the “axolotl” module.
-To favor selective updates, DitHub performs a fetch operation (Fig.
-2) to retrieve any previously stored version of the specialized module, such as $A_{\text{dog}}^{\text{old}}$.
-If found, this module is merged (merge in Fig.
-2) with the current task’s warmup matrix $A_{\text{wu}}$.
-For instance, considering the class “dog”, the merge operation follows a weighted strategy:
-$A_{\text{dog}}^{\text{cur}}=(1-\lambda_{A})A_{\text{dog}}^{\text{old}}+\lambda_{A}A_{\text{wu}}$ (3)
-Here, $\lambda_{A}$ is set to a low value, prioritizing consolidated knowledge over new insights from the warmup phase.
-Details on the impact of $\lambda_{A}$ are in Sec.
-5.3.
-Crucially, our fetching process creates an on-the-fly initialization that blends historical class-specific knowledge with new domain information.
 Algorithm 1 DitHub Training
-Input: $T$ disjoint tasks each containing $|\mathcal{C}_{t}|$ classes; library $\mathcal{A}\leftarrow\{\}$ of class experts; global $B$.
+Input: $T$ tasks each containing $|\mathcal{C}_{t}|$ classes; library $\mathcal{A}\leftarrow\{\}$ of class experts; global $B$.
 for task $t\in\{1,\dots,T\}$ do
@@
          if $A_{c}\in\mathcal{A}$ then $\triangleright$  Fetch
-              $A_{c}^{\text{cur}}\leftarrow$ use Equation 3 $\triangleright$  Merge
+              $A_{c}^{\text{cur}}\leftarrow$ use Eq. 2 $\triangleright$  Merge
          else
@@
          $\mathcal{A}\leftarrow\mathcal{A}\cup A_{c}$
-     $B\leftarrow$ use Equation 4
+     $B\leftarrow$ use Eq. 3
+Figure 3: Memory requirements as the number of classes grows.
+To favor selective updates, DitHub performs a fetch operation (Fig.
+2), which simply retrieves the class-specific module from previous tasks if the class has already been seen — followed by the merge (Fig.
+2) .
+For instance, considering the class “dog”, the previously stored expert module $A_{\text{dog}}^{\text{old}}$ is fetched and merged with the current task’s warmup matrix $A_{\text{wu}}$ following a weighted strategy:
+$A_{\text{dog}}^{\text{cur}}=(1-\lambda_{A})A_{\text{dog}}^{\text{old}}+\lambda_{A}A_{\text{wu}}$ (2)
+Here, $\lambda_{A}$ is set to a low value, prioritizing consolidated knowledge over new insights from the warmup phase.
+Details on the impact of $\lambda_{A}$ are in Sec.
+5.2.
+Crucially, our merging process creates an on-the-fly initialization that blends historical class-specific knowledge with new domain information, resulting in $A_{\text{dog}}^{\text{cur}}$, which serves as the starting point for the specialization phase.
 4.2 Bridging Modularity with Memory Efficiency
-Figure 3: Memory requirements as the number of classes grows.
 To keep our library memory-efficient, we maintain a single shared $B$ matrix instead of a distinct $B$ for each class.
@@
 A.2, we present DitHub’s results when $B$ is not shared..
-Indeed, considering DitHub “$B$ shared” (line [Uncaptioned image]), the memory footprint of our approach is consistently lower than the storage needed for the Grounding DINO encoder model alone (line [Uncaptioned image]).
+Additionally, considering DitHub “$B$ shared” (line [Uncaptioned image]), the memory footprint of our approach is consistently lower than the storage needed for the Grounding DINO encoder alone (line [Uncaptioned image]).
 This makes DitHub’s memory demands manageable for practical applications.
-Also, the memory efficiency is further optimized by fine-tuning only a limited number of layers (see App.
-A).
+Moreover, memory efficiency is further optimized by fine-tuning a targeted portion of the base architecture (Sec.
+A.1).
 Due to the shared design, the matrix $B$ is more susceptible to interference and catastrophic forgetting [30].
-Consequently, we implement a tailored mechanism that ensures smooth updates as tasks progress.
+Consequently, we implement a mechanism that ensures smooth updates as tasks progress.
 Specifically, at the end of each task $t$, we update the matrix $B$ by merging the previous checkpoint, $B_{t-1}$, with the weights learned after the current task, $B^{\text{opt}}$.
@@
 Unlike Eq.
-3, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state.
-This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, which will be further explained in Sec.
-5.3.
-4.3 Detection with Class-Specific Modules
-With our library of efficient, class-specific modules in place for adapting the Vision-Language backbone, two disjoint inference routes are available for users.
-Given a textual query specifying the classes of interest, our model tokenizes the prompt and searches token by token to determine if corresponding expert modules are present in the library.
-If there are correspondences, those relevant modules are fused using a simple average to create a single composite adaptation module.
-Conversely, for requests targeting individual classes where fine-grained adaptation is desired, the corresponding module can be directly selected and deployed.
+2, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state.
+This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, further detailed in Sec.
+5.2.
+4.3 Inference with DitHub
+With our library of efficient, class-specific modules in place for adapting the Vision-Language detector, DitHub enables two complementary inference strategies.
+For requests targeting individual classes where fine-grained adaptation is needed, the corresponding expert module can be directly selected and used to detect that class.
+Conversely, when evaluating on a textual prompt that includes multiple classes, we retrieve the corresponding class-specific modules from the library and merge them by computing an average of their weights to create a single composite adaptation module.
+This unified module is then used in a single forward pass to perform detection over all target classes.
 In our experimental studies, we present results and discussions that address both scenarios.
 5 Experimental Studies
-This section evaluates DitHub in two parts.
-The first assesses its adaptability to multiple tasks using the Incremental Vision-Language Object Detection (IVLOD) protocol [2] in Sec.
-5.1.
-The second, in Sec.
-5.2, examines the method’s specialization capabilities.
-The section ends with ablation studies in Sec.
-5.3.
 5.1 Incremental Vision-Language Object Detection
+Evaluation settings.
+Following [4], we employ ODinW-13 [20] to assess the model ability to learn new tasks while mitigating forgetting.
+ODinW-13 includes $13$ sub-datasets, spanning from the traditional Pascal VOC [8] to more challenging domains with significant distribution shifts, such as thermal imagery.
+Also, the IVLOD setting requires the Open-Vocabulary detector to retain strong zero-shot capabilities after fine-tuning, which are evaluated on MS COCO [22].
+One issue of ODinW-13 is that most classes are confined to a single task among the $13$.
+However, DitHub is conceived to handle recurring classes across domains, necessitating updates to their respective modules.
+To better assess incremental detection and module reuse, we introduce ODinW-O (Overlapped), a subset of ODinW-35 that is designed to reflect realistic scenarios where classes can reoccur across tasks over time.
+Further details are provided in App.
+C.
+For comparisons, we employ mean Average Precision (mAP) and compare DitHub against six state-of-the-art methods.
+CL-DETR [25] and OW-DETR [11] are explicitly designed for Incremental Learning, while TFA [48] specializes in few-shot adaptation.
+AT [12] has been extended to both incremental and few-shot learning, whereas iDETR [6] natively integrates both aspects.
+Finally, ZiRa [4], our primary competitor, is specifically tailored for IVLOD.
+Implementation details.
+We focus on Grounding DINO and adapt only its encoder network, as further validated in Sec.
+A.1.
+The merging coefficient $\lambda$ for the $A$ matrix is set to $0.3$ for ODinW-13 and $0.1$ for ODinW-O; for the matrix $B$, we set $\lambda=0.7$.
+The choice of $\lambda$ and its underlying rationale are further discussed in Sec.
+5.2.
+During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — see App.
+E for further hyperparameter details.
+Performance on ODinW-13.
+Tab.
+1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO).
+We evaluate performance in both the full-data setting and few-shot scenarios with $1$, $5$, and $10$ shots per class.
+All baseline results, except those of Grounding DINO, ZiRa, and our method, are taken directly from [4].
+Each number represents the average of three runs with different random seeds, with standard deviations provided in Sec.
+D.1.
+Table 1: Comparison of mAP values across ODinW-13 tasks. Avg represents the average across the $13$ tasks, while zero-shot performance is reported in the ZCOCO column. Bold indicates best results.
 Shots Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
@@
 Full TFA $30.97$ $47.93$ $23.80$ $30.65$ $67.21$ $61.77$ $30.52$ $50.23$ $47.73$ $60.91$ $29.25$ $61.72$ $31.42$ $66.23$ $61.61$
-iDETR 37.32 58.71 32.64 46.65 70.99 68.56 55.32 58.88 64.48 71.01 50.33 63.30 39.19 77.12 64.80
-AT 42.30 51.14 23.62 39.90 72.32 65.51 31.47 50.48 60.51 66.07 39.09 53.50 34.04 68.07 60.23
+iDETR $37.32$ $58.71$ $32.64$ $46.65$ $70.99$ $68.56$ $\boldsymbol{55.32}$ $58.88$ $64.48$ $71.01$ $50.33$ $63.30$ $39.19$ $77.12$ $64.80$
+AT $42.30$ $51.14$ $23.62$ $39.90$ $\boldsymbol{72.32}$ $65.51$ $31.47$ $50.48$ $60.51$ $66.07$ $39.09$ $53.50$ $34.04$ $68.07$ $60.23$
 OW-DETR $31.22$ $55.58$ $28.46$ $43.78$ $70.54$ $67.78$ $43.84$ $56.75$ $63.13$ $69.51$ $45.16$ $58.99$ $36.99$ $74.42$ $63.20$
-CL-DETR 32.15 57.26 29.35 45.15 71.94 69.90 45.21 58.52 65.10 71.68 46.58 60.83 38.14 76.74 65.18
+CL-DETR $32.15$ $57.26$ $29.35$ $45.15$ $71.94$ $\boldsymbol{69.90}$ $45.21$ $58.52$ $65.10$ $\boldsymbol{71.68}$ $46.58$ $60.83$ $38.14$ $76.74$ $65.18$
 ZiRA $46.26$ $57.98$ $31.76$ $47.35$ $71.77$ $64.74$ $46.53$ $62.66$ $66.39$ $71.00$ $48.48$ $63.03$ $41.44$ $76.13$ $62.44$
-DitHub 47.01 62.19 34.62 50.65 70.46 68.56 49.28 65.57 69.58 71.10 56.65 70.88 52.82 79.30 68.18
-Table 1: Comparison of mAP values across ODinW-13 tasks.
-Avg represents the average across the $13$ tasks, while zero-shot performance is reported in the ZCOCO column.
-Best results are highlighted in bold.
-Evaluation settings.
-Following [2], we employ ODinW-13 [17] to assess the model ability to learn new tasks while mitigating forgetting of previously seen data.
-ODinW-13 consists of $13$ sub-datasets, spanning from the traditional Pascal VOC [6] to more challenging domains with significant distribution shifts, such as thermal imagery.
-Also, the IVLOD setting mandates that the Open-Vocabulary detector maintains robust zero-shot capabilities after fine-tuning.
-These capabilities are tested on MS COCO [19].
-One issue of ODinW-13 is that most classes are confined to a single task among the $13$.
-However, DitHub is conceived to handle classes that reoccur across different domains, necessitating updates to their respective modules.
-To provide an additional perspective on incremental object detection and model reuse, we introduce a variant of ODinW-13, referred to as ODinW-O.
-Starting from ODinW-35, we construct ODinW-O by retaining only the classes that are shared across tasks while discarding those that appear in only one task.
-Further details are provided in App.
-B.
-For comparisons, we employ mean Average Precision (mAP) and compare DitHub against six state-of-the-art methods.
-CL-DETR [22] and OW-DETR [9] are explicitly designed for incremental learning, while TFA [43] specializes in few-shot adaptation.
-AT [10] has been extended to both incremental and few-shot learning, whereas iDETR [4] integrates both aspects.
-Finally, ZiRa [2], our primary competitor, is specifically tailored for IVLOD.
-Implementation details.
-We focus on Grounding DINO and adapt only its encoder network, as further validated in Sec.
-A.1.
-The merging coefficient $\lambda$ for the $A$ matrix is set to $0.3$ for ODinW-13 and $0.1$ for ODinW-O; for the matrix $B$, we set $\lambda=0.7$.
-The choice of $\lambda$ and its underlying rationale are further discussed in Sec.
-5.3.
-During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — App.
-C for further hyperparameter details.
-Performance on ODinW-13.
-Table 1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO).
-We evaluate performance in both the standard setting with the full dataset and the few-shot scenarios, considering 1, 5, and 10 shots per class.
-All baseline results, except those of Grounding DINO, ZiRa, and our method, are taken directly from [2].
-Each number represents the average of three runs with different random seeds, with standard deviations provided in Sec.
-D.1.
-In the full-shot setting, DitHub outperforms its main competitor, ZiRa, by a substantial + 4.21 mAP points and achieves the best results in nine out of thirteen tasks.
-Notably, on ZCOCO, our approach outperforms ZiRa by a mAP gain of + 0.75, setting a new state-of-the-art in both incremental and zero-shot retention capabilities.
-Although DitHub is not explicitly designed for data-scarce conditions, DitHub achieves state-of-the-art performance even in the few-shot setting.
-In particular, in the five-shot and ten-shot settings, DitHub emerges as the best-performing approach, with the ten-shot scenario showing a slightly larger improvement over ZiRa (+ 1.23 mAP).
-Other baselines struggle to remain competitive in this regime.
-Our method also performs well in the one-shot setting, where iDETR achieves the highest average performance, followed closely by DitHub and ZiRa.
-These findings — optimal performance in incremental learning, effective retention of zero-shot capabilities, and robustness in the low-data regime — clearly highlight the advantages of our modular approach, with the advanced update system that will be dissected in the analyses below.
-ZCOCO Avg Ae Hw Pv Sd Th Ve
+DitHub $\boldsymbol{47.01}$ $\boldsymbol{62.19}$ $\boldsymbol{34.62}$ $\boldsymbol{50.65}$ $70.46$ $68.56$ $49.28$ $\boldsymbol{65.57}$ $\boldsymbol{69.58}$ $71.10$ $\boldsymbol{56.65}$ $\boldsymbol{70.88}$ $\boldsymbol{52.82}$ $\boldsymbol{79.30}$ $\boldsymbol{68.18}$
+In the full-shot setting, DitHub outperforms ZiRa by a substantial $\boldsymbol{\texttt{+}4.21}$ mAP and achieves the best results in nine out of thirteen tasks.
+Notably, on ZCOCO, it surpasses ZiRa by a $\boldsymbol{\texttt{+}0.75}$ mAP, setting a new state-of-the-art in both incremental and zero-shot retention.
+While not explicitly designed for data-scarce conditions, DitHub also excels in the few-shot setting.
+In particular, in the five-shot and ten-shot settings, DitHub emerges as the best-performing approach, with the ten-shot scenario showing a slightly larger improvement over ZiRa ($\boldsymbol{\texttt{+}1.23}$ mAP).
+Our method also performs well in the one-shot setting, where iDETR leads, followed closely by DitHub and ZiRa.
+These findings – optimal performance in Incremental Learning, effective retention of zero-shot capabilities, and robustness in the low-data regime – clearly highlight the advantages of our modular approach.
+Performance on ODinW-O.
+The ODinW-O scenario emphasizes that frequently encountered classes may reappear in subsequent tasks or at later times, such as the class “person” in common industrial and smart-city scenarios.
+As detailed in Sec.
+C.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks.
+In this benchmark, we primarily compare our approach with ZiRa, as it is the most competitive method in the standard IVLOD setting, excluding ours.
+Table 2: mAP values on ODinW-O. Best in bold.
+Method ZCOCO Avg Ae Hw Pv Sd Th Ve
 G-Dino $47.41$ $53.15$ $45.12$ $67.54$ $58.11$ $25.84$ $70.40$ $51.87$
-ZiRa 44.43 57.63 39.92 68.00 64.90 46.26 77.26 49.47
-DitHub 46.51 62.38 53.35 71.07 71.01 41.75 80.21 56.90
-Table 2: mAP values on ODinW-O. Best in bold.
+ZiRa $44.43$ $57.63$ $39.92$ $68.00$ $64.90$ $\boldsymbol{46.26}$ $77.26$ $49.47$
+DitHub $\boldsymbol{46.51}$ $\boldsymbol{62.38}$ $\boldsymbol{53.35}$ $\boldsymbol{71.07}$ $\boldsymbol{71.01}$ $41.75$ $\boldsymbol{80.21}$ $\boldsymbol{56.90}$
+The results in Tab.
+2 indicate that our approach secures a substantial improvement of $\boldsymbol{\texttt{+}4.75}$ in mAP and a $\boldsymbol{\texttt{+}2.08}$ mAP gain on ZCOCO over ZiRa.
+We attribute the success of our method to its class-oriented modular design, which enables selective updates to recurring concepts.
+This design prevents unintentional overwriting of knowledge across classes, tasks, and domains.
+Figure 4: Normalized Average mAP as tasks progress.
+To assess performance degradation, we normalize the mAP for each task by dividing the current value by the original mAP recorded at the end of that task, i.e., if the curve $\approx 0.90$, the model retains $90\%$ of its original performance.
+Performance analysis as tasks progress.
+In Fig.
+4, we present an incremental analysis comparing the performance of DitHub and its strongest competitor, ZiRa, after each task.
+The evaluation, conducted on both ODinW-13 (left) and ODinW-O (right), allows us to assess the retention of both zero-shot and in-domain capabilities throughout the task sequence, providing a direct measure of their resistance to forgetting (further discussed in App.
+B).
+On ZCOCO, DitHub exhibits stable performance across tasks, with a negligible loss in zero-shot capabilities — unlike ZiRa, which shows a more pronounced decline.
+When evaluating performance across the sequence of downstream tasks, both methods demonstrate a decreasing trend, indicating that forgetting remains a significant challenge in Incremental Object Detection.
+Nevertheless, as more tasks are introduced, the performance gap widens, with DitHub consistently outperforming its competitor.
+This highlights the effectiveness of our modular design, which enables scalable knowledge organization and adaptation as the number of tasks grows.
+Moreover, the gap is still present in the ODinW-O setting introduced in this work, which models the realistic scenario where subsequent tasks may share classes.
+This setting, being inherently more susceptible to forgetting, further underscores the advantages of our approach.
+5.2 Model Analysis and Ablation Studies
+We examine the properties of our specialization framework and the benefits of composing class-specific experts.
+To this aim, we introduce an ablative approach, Experts-not-Experts (EnE), which removes class specialization.
+EnE is identical to DitHub, except during training, where it omits the stochastic strategy from Sec.
+4.1.
+Instead, given an image from the current task, EnE performs the training step on one module chosen at random among all task classes.
+As a result, modules no longer specialize in a class but instead learn from the whole task data, including images without their designated class.
+Consequently, they become generalist learners rather than class-specific experts.
+In Tab.
+3, we compare DitHub with its non-specialized variant, i.e., EnE.
+Notice that, at inference, EnE acts as an ensemble of class-agnostic modules.
+In contrast, DitHub employs class-specialized counterparts with an expertise on data containing the same class.
+This targeted specialization yields a $\boldsymbol{\texttt{+}1.23}$ mAP improvement on average, with additional gains also on the ZCOCO metric.
+These results affirm that a system of modular experts, each with distinct knowledge, offers a promising approach for systems requiring continuous and selective adaptation.
+Table 3: Experts-not-Experts (EnE). Training by selecting a random expert.
 Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
@@
 DitHub $47.01$ $62.19$ $34.62$ $50.65$ $70.46$ $68.56$ $49.28$ $65.57$ $69.58$ $71.10$ $56.65$ $70.88$ $52.82$ $79.30$ $68.18$
-Table 3: Experts not Experts (EneE). Training by selecting a random expert.
-Performance on ODinW-O.
-The ODinW-O scenario emphasizes that frequently encountered classes may reappear in subsequent tasks or at later times, such as the class “person” in common industrial and smart-city scenarios.
-As detailed in Sec.
-B.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks.
-In this benchmark, we primarily compare our approach with ZiRa, as it is the most competitive method in the standard IVLOD setting, excluding ours.
-The results in Tab.
-2 indicate that our approach secures a substantial improvement of + 4.75 in mAP and a + 2.08 mAP gain on ZCOCO over ZiRa.
-We attribute the success of our method to its class-oriented modular design, which enables selective updates to recurring concepts.
-This design effectively prevents the model from inadvertently overwriting crucial knowledge associated with other classes, tasks, or domains.
-5.2 Evaluation of Specialized Detection Modules
-Here, we delve into the properties of our specialization framework and the benefits of composing class-specific experts.
-To this aim, we introduce an ablative approach called Experts-not-Experts (EnE), which emphasizes the absence of class specialization in our experts.
-EnE is identical to DitHub, except during training, where it does not employ the stochastic strategy introduced in Sec.
-4.1 and designed to specialize the experts.
-Instead, given an image from the current task, the EnE approach performs the training step on one module chosen at random among the classes of the entire task.
-As a result, modules no longer specialize in a class but instead learn from the whole task data, including images without their designated class.
-Consequently, they become general task learners rather than class-specific experts.
-Figure 4: Specialization capabilities for rare classes of ODinW-13.
-In Tab.
-3, we offer a comparative analysis between our DitHub and its non-specialized variant, i.e., EnE.
-Notice that, at inference time, EnE acts as an ensemble of modules, each trained on the same data but with a different random ordering.
-In contrast, DitHub employs class-specialized modules with an expertise on data containing the same class.
-This targeted specialization improves performance, as shown in Tab.
-3, where per-class specialization results in a + 1.23 mAP improvement on average.
-Additionally, albeit to a lesser extent, the ZCOCO metric still benefits from an improvement.
-These results affirm that a system composed of modular experts, each a custodian of distinct knowledge, offers a promising approach for systems requiring continuous and selective adaptation.
+Figure 5: Results on rare classes of ODinW-13.
+Figure 6: Showing unlearning skills of DitHub.
 Evaluation on rare classes.
-As discussed in Section 4.3, DitHub facilitates fine-grained control during inference, enabling users to tailor the system to their specific use case and deploy the most appropriate modules.
-This includes the option to use a single module specialized on a target class.
-To assess single-module effectiveness, we conduct a targeted evaluation on rare classes from ODinW-13, activating only the corresponding target class module.
+As discussed in Sec.
+4.3, DitHub offers fine-grained control at inference, enabling users to tailor the system to their specific use case — including the option to use a single module specialized for a target class.
+To assess this, we conduct a targeted evaluation on rare classes from ODinW-13, activating only the corresponding target class module.
 Results in Fig.
-4 show that DitHub consistently achieves the best results across rare classes, highlighting the benefits of class specialization.
-EnE and DitHub outperform the zero-shot Grounding DINO baseline, though EnE underperforms for the Starfish category.
-Figure 5: Showing the natural unlearning capabilities of DitHub.
+6 show that DitHub achieves the best results on rare classes, highlighting the benefits of class specialization.
 Specialized modules allows training-free unlearning.
-The advantages of DitHub extend beyond additive incremental knowledge compositionality.
-Indeed, we can show that expert modules can be useful to undermine or remove some specific knowledge or content from a model, by simply subtracting the corresponding weights from the base model.
-This operation, which resembles the concept of zero-training unlearning, has been explored by the authors of [13, 28] in the context of image and text categorization tasks; to the best of our knowledge, we are the first to explore its application within the context of object detection.
-Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class module from the pre-trained model: $\mathbf{W}^{\prime}=\mathbf{W}-\alpha BA_{c}$, where we set $\alpha=0.3$.
-As a result, we provide a qualitative visualization shown in Fig.
-5: on the left, the zero-shot model correctly recognizes common classes, while on the right, applying the negative module of a specific class reduces its mAP, effectively unlearning it and leading to false negatives.
-Additionally, we provide a quantitative evaluation of the unlearning process in Sec.
+DitHub’s advantages go beyond additive Incremental Learning.
+Expert modules can also be used to remove specific knowledge by simply subtracting the corresponding weights from the base model.
+This operation, akin to zero-training unlearning, has been studied in image and text categorization [15, 33]; to the best of our knowledge, we are the first to explore it within the context of Object Detection.
+For example, this enables the model to unlearn the detection of sensitive object categories – such as faces, license plates, or medical equipment – when required to comply with privacy regulations or data retention policies.
+Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class $c$ module from the pre-trained model: $W^{\prime}=W-\alpha BA_{c}$, whit $\alpha=0.3$.
+Fig.
+6 illustrates this qualitatively: on the left, the zero-shot model correctly detects common classes, while on the right, subtracting a class module reduces its mAP, effectively unlearning it and leading to false negatives.
+A quantitative analysis is provided in Sec.
 D.2.
-5.3 Ablation Studies
-Figure 6: Sensitivity analysis on $\lambda_{A}$ and $\lambda_{B}$. Performance measured as harmonic mean of Avg mAP (across tasks) and ZCOCO.
-Sensitivity Analysis of $\lambda.$ Analyzing the variation of the merging coefficients $\lambda_{A}$ and $\lambda_{B}$ offers valuable insights into our methodology.
-To show that, we define an aggregate metric that is the harmonic mean between DitHub’s full-shot performance on ZCOCO and Avg, ensuring a balanced evaluation of specialization (Avg) and zero-shot generalization (ZCOCO).
-Figure 6 presents the performance trends for ODinW-13 (top) and ODinW-O (bottom) while varying $\lambda_{A}$ and $\lambda_{B}$ in increments of $0.2$.
+Sensitivity Analysis of $\boldsymbol{\lambda}$.
+Varying the merging coefficients $\lambda_{A}$ and $\lambda_{B}$ offers valuable insights into our method.
+To this end, we define an aggregate metric as the harmonic mean of DitHub’s full-shot performance, balancing specialization (Avg) and zero-shot generalization (ZCOCO).
+Fig.
+8 shows performance trends for ODinW-13 (top) and ODinW-O (bottom) as $\lambda_{A}$ and $\lambda_{B}$ in $0.2$ increments.
 Considering $\lambda_{A}$, its role is significant while merging the expert $A$ matrices (see Eq.
-3).
+2).
 From Fig.
-6, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective.
+8, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective.
 Interestingly, while $\lambda_{A}=0.3$ performs best for ODinW-13, ODinW-O exhibits a monotonic decline as $\lambda_{A}$ increases.
-This contrast arises from dataset structures: ODinW-O consists solely of overlapping classes across tasks, where retaining past expertise is critical, whereas ODinW-13 features diverse class distributions per task, requiring greater integration of domain-specific warmup knowledge.
+The contrast reflects dataset structure: ODinW-O includes overlapping classes, emphasizing knowledge retention, whereas ODinW-13 has varied class distributions, demanding stronger integration of domain-specific warmup knowledge.
 For the incremental adaptation of the shared $B$ matrix, a different trend emerges.
@@
 This suggests that prioritizing the knowledge in the $B^{\text{opt}}$ matrix from the current task, as shown in Eq.
-4, is crucial.
+3, is crucial.
 Our intuition, supported by [29], is that the latest $B^{\text{opt}}$ matrix integrates knowledge from both earlier and recent tasks, making it more valuable than $B_{t-1}$, which is trained solely on earlier tasks.
+Figure 7: Sensitivity analysis on $\lambda_{A}$ and $\lambda_{B}$. Performance measured as harmonic mean of Avg and ZCOCO mAP.
 Figure 8: Ablative study on the DitHub components. Results are expressed as average performance across the tasks of ODinW-13.
-On Warmup and Shared B.
-To evaluate the impact of each component in DitHub, we progressively integrate them and present the results in Fig.
-7.
-We start with the zero-shot performance of Grounding DINO (GD) as a reference, followed by a baseline fine-tuning procedure (Base in Fig.
-7), where an $A$ matrix is trained for each class without merging strategies, omitting the warmup phase and using a shared $B$ matrix.
-This method leads to complete forgetting as new tasks are introduced, but still outperforms Grounding DINO, which struggles with pretraining knowledge alone.
-Building on this baseline, we first incorporate Eq.
-4, but see no significant performance gain.
-This is expected, as the shared $B$ matrix only represents a tiny portion of the total parameters.
-Without class-specific focus, Eq.
-4 alone does not enhance retention.
-However, adding the warmup phase (+wu) provides a clear improvement, as it stabilizes training and helps rare classes by offering a robust initialization.
-When we further include Eq.
-4 in the warmup-enhanced fine-tuning strategy, a modest improvement is observed.
-However, the most substantial performance gain occurs when Eq.
-3 is also integrated into the methodology (DitHub in Fig.
-7).
-This highlights the importance of preserving past domain knowledge within class-specific modules, ensuring a more stable and effective incremental learning process.
-Notably, the combination of all three key components — the warmup phase for reliable initialization, the merging strategy for $A$ matrices to retain class-specific knowledge across domains, and the shared $B$ matrix approach for efficiency — results in a significant + 6.22 improvement over the baseline where none of these strategies are applied.
+On Warmup and Shared $\boldsymbol{B}$.
+To evaluate the impact of DitHub’s components, we progressively integrate them and present results in Fig.
+8.
+We establish the zero-shot performance of Grounding DINO (GD) as a reference.
+Our baseline fine-tuning procedure (Base) trains class-specific $A$ matrices and the shared $B$ matrix independently for each new task, omitting both parameter merging and the warmup phase.
+While this baseline suffers from complete catastrophic forgetting, it still demonstrates improved performance compared to the zero-shot Grounding DINO.
+Building on this baseline, we first add the warmup phase (+wu), which yields a notable improvement by stabilizing training and facilitating the learning of rare classes through robust initialization.
+Subsequently, incorporating Eq.
+3 into the warmup-augmented strategy results in a modest improvement — this is expected, as the shared $B$ matrix constitutes only a small fraction of the total parameters.
+Crucially, the most substantial performance gain is achieved upon integrating Eq.
+2 (DitHub in Fig.
+8), highlighting the importance of retaining domain-specific knowledge within class-specific modules to support stable and effective Incremental Learning.
+Collectively, the synergistic combination of these three key components – the warmup phase for robust initialization, the merging strategy for $A$ matrices to retain class-specific knowledge across domains, and the shared $B$ matrix for efficiency – yields a substantial $\boldsymbol{\texttt{+}6.13}$ mAP improvement compared to the baseline model.
 6 Conclusions
@@
 The incremental nature of this domain allowed us to explore the compositional capabilities of efficient modules, marking the first such investigation in Object Detection.
-Looking ahead, we aim to further leverage DitHub’s modular library management to deepen our understanding of knowledge transfer between efficient modules for object detection.
+In future works, we aim to further leverage modular library management to deepen our understanding of knowledge transfer between efficient modules for Object Detection.
 For instance, if our library contains an “axolotl” module trained on real-world images and we possess domain knowledge for the cartoon image domain, can we synthesize a module specialized in detecting cartoon “axolotls” through knowledge transfer?
 DitHub will provide a fork of this manuscript to explore such questions.
-Supplementary Material Overview
-This supplementary material complements the main manuscript by providing additional analyses, insights, and results on DitHub.
-Specifically, we delve deeper into efficiency considerations of our adaptation approach (App.
-A), highlighting blockwise adaptation and LoRA specializations.
-We introduce the novel ODinW-O benchmark for incremental learning with overlapping classes and domain shifts (App.
-B), and present insights on unlearning class-specific knowledge (Sec.
-D.2).
-Additionally, we clarify critical implementation specifics (App.
-C), present extensive robustness checks (Sec.
-D.1), and showcase qualitative examples (App.
-E) to illustrate the practical benefits of our methodology.
-Appendix A Further Insights on Efficiency
-A.1 Blockwise Adaptation Analysis
-While Grounding DINO consists of four main components — an image backbone for visual encoding, a textual backbone for language understanding, and an encoder-decoder pair for detection — we focus our adaptation on the latter two.
-This choice is motivated by the need for a more detection-centric adaptation and efficiency considerations, as the textual backbone, being a BERT Base model [3], accounts for more than half of Grounding DINO ’s parameters.
-In Table B, we present results for adapting only the encoder (Figure A), only the decoder (Figure A), and both encoder and decoder (Figure A).
-DitHub employs the first approach — adapting only the encoder — as it delivers the best performance on both ZCOCO and Avg metrics.
-A.2 Employing Full LoRA for Class-Specificity
-While our methodology leverages a shared and unique $B$ matrix for all tasks, Tab.
-A presents results obtained using class-specific pairs.
-In this variant of DitHub, both the $A$ and $B$ low-rank matrices of LoRA undergo the same specialized optimization process: an initial warmup phase followed by dedicated specialization.
-In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Section 4.1.
-Applying our full set of strategies to both matrices yields a + 1.03 improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Fig.
-3).
-Prioritizing efficiency, we opt for a lightweight implementation that trades a marginal performance gain for significantly reduced memory requirements.
-Notably, ZCOCO results favor our efficient approach, as a single, unspecialized $B$ matrix remains more aligned with the pretraining distribution, thereby mitigating degradation in zero-shot capabilities.
-Method ZCOCO Avg
-DitHub 47.01 62.19
-DitHub (LoRA full) 46.66 63.22
-Table A: LoRA full results on ODinW-13, with DitHub results taken from Table 1 for reference.
-(a) Encoder
-(b) Decoder
-(c) Decoder/Encoder
-Figure A: Blockwise analysis of adaptation strategies: (a) Adaptation of the encoder only (our chosen approach), (b) Adaptation of the decoder only and (c) Joint adaptation of both the decoder and encoder.
+Appendix Overview
+This Appendix provides supplementary information regarding the main manuscript.
+  • App. A — Further Insights on Efficiency: Optimizes DitHub’s performance when adapting Grounding DINO.
+    • Sec. A.1 — Blockwise Adaptation Analysis: Justifies adapting only the encoder for DitHub based on empirical metrics.
+    • Sec. A.2 — Employing Full LoRA for Class-Specificity: Explores class-specific LoRA, favoring a parameter-efficient shared $B$ matrix.
+  • App. B — Further Insights on Forgetting: Analyzes DitHub’s anti-forgetting capabilities.
+    • Sec. B.1 — Forgetting on ODinW-13: Presents anti-forgetting results for ODinW-13.
+    • Sec. B.2 — Forgetting on ODinW-O: Demonstrates robust anti-forgetting on the ODinW-O benchmark.
+  • App. C — Additional Benchmark Details: Provides comprehensive dataset information.
+    • Sec. C.1 — ODinW-13 and MS COCO: Describes ODinW-13 for specialization and MS COCO (ZCOCO) for zero-shot evaluation.
+    • Sec. C.2 — ODinW-O: Introduces the ODinW-O benchmark for Incremental Learning and domain shifts.
+  • App. D — Additional Results: Presents complementary experimental information.
+    • Sec. D.1 — Standard Deviations: Reports standard deviations for reproduced methods from Tab. 1 and Tab. 2.
+    • Sec. D.2 — Unlearning Additionals: Illustrates DitHub’s modular capability for selective knowledge removal (Fig. C).
+  • App. E — Implementation Details: Outlines experimental setup, including hardware, model configurations, and optimization specifics.
+  • App. F — Limitations and Broader Impact: Discusses current limitations and broader societal implications.
+  • App. G — Qualitative Results: Provides visual comparisons (Fig. D) demonstrating DitHub’s improved detection and robustness.
+Table A: Performance comparison across datasets when adapting only the decoder (Only Dec), both encoder and decoder (Dec/Enc), and encoder only (DitHub).
 Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
@@
 Dec/Enc $46.55$ $59.92$ $32.84$ $43.16$ $71.47$ $67.44$ $49.78$ $62.84$ $70.32$ $70.83$ $49.24$ $70.89$ $46.12$ $77.05$ $67.48$
-Table B: Performance comparison across datasets when adapting only the decoder (”Only Dec”), both encoder and decoder (”Dec/Enc”), and encoder only (DitHub).
-Appendix B Additional Benchmark Details
+(a) Encoder
+(b) Decoder
+(c) Decoder/Encoder
+Figure A: Blockwise analysis of adaptation strategies: (a) encoder-only adaptation (our selected method), (b) decoder-only adaptation, and (c) joint encoder-decoder adaptation.
+Appendix A Further Insights on Efficiency
+A.1 Blockwise Adaptation Analysis
+While Grounding DINO consists of four main components – an image backbone for visual encoding, a textual backbone for language understanding, and an encoder-decoder pair for detection – we focus our adaptation on the latter two.
+This choice is motivated by the need for a more detection-centric adaptation and efficiency considerations, as the textual backbone, being a BERT Base model [5], accounts for more than half of Grounding DINO ’s parameters.
+In Tab.
+A, we present results for adapting only the encoder (Fig.
+A), only the decoder (Fig.
+A), and both encoder and decoder (Fig.
+A).
+DitHub employs the first approach – adapting only the encoder – as it delivers the best performance on both ZCOCO and Avg metrics.
+A.2 Employing Full LoRA for Class-Specificity
+Table B: LoRA full results on ODinW-13, with DitHub results taken from Tab. 1 for reference.
+Method ZCOCO Avg
+DitHub $47.01$ $62.19$
+DitHub (LoRA full) $46.66$ $63.22$
+While our methodology leverages a shared and unique $B$ matrix for all tasks, Tab.
+B presents results obtained using class-specific pairs.
+In this variant of DitHub, both the $A$ and $B$ low-rank matrices of LoRA undergo the same specialized optimization process: an initial warmup phase followed by dedicated specialization.
+In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Sec.
+4.1.
+Applying our full set of strategies to both matrices yields a $\boldsymbol{\texttt{+}1.03}$ mAP improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Fig.
+3).
+Prioritizing efficiency, we opt for a lightweight implementation that trades a marginal performance gain for significantly reduced memory requirements.
+Notably, ZCOCO results favor our efficient approach, as a single, unspecialized $B$ matrix remains more aligned with the pretraining distribution, thereby mitigating degradation in zero-shot capabilities.
+Appendix B Further Insights on Forgetting
+This section presents an additional analysis of the anti-forgetting properties of DitHub, complementing the study in Sec.
+5.2 by offering an alternative perspective.
+We report the forgetting metric for both ODinW-13 and ODinW-O, computed for DitHub and its primary competitor, ZiRa [4].
+This metric is calculated using the canonical definition [3, 1]: the difference in performance on a specific task after learning the entire sequence of tasks compared to the original performance on that task immediately after its initial learning.
+B.1 Forgetting on ODinW-13
+Table C: Forgetting metrics in terms of mAP for the ODinW-13 dataset. The numbers above each task indicate the order in which they were processed. Bold indicates best results.
+$11$ $10$ $2$ $13$ $12$ $1$ $7$ $6$ $3$ $4$ $9$ $8$ $5$
+Method Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
+ZiRa -$\texttt{-}6.66$$\texttt{-}4.31$$\texttt{-}4.82$$\boldsymbol{\texttt{-}1.80}$ $0.00$$\texttt{-}2.09$$\texttt{-}8.71$$\texttt{-}5.80$$\texttt{-}7.76$$\texttt{-}12.63$$\texttt{-}19.19$$\boldsymbol{\texttt{-}1.82}$$\texttt{-}9.85$$\texttt{-}7.77$
+DitHub -$\boldsymbol{\texttt{-}1.73}$$\boldsymbol{\texttt{-}1.56}$$\boldsymbol{\texttt{-}1.69}$$\texttt{-}2.34$ $0.00$$\boldsymbol{\texttt{-}1.61}$ $\boldsymbol{~{}~{}3.62}$$\boldsymbol{\texttt{-}1.91}$ $\boldsymbol{~{}~{}1.54}$$\boldsymbol{\texttt{-}6.94}$$\boldsymbol{\texttt{-}7.27}$$\texttt{-}2.11$ $\boldsymbol{~{}~{}0.10}$$\boldsymbol{\texttt{-}2.34}$
+Tab.
+C presents the forgetting in terms of mAP points for the $13$ tasks of ODinW-13, with the average across all tasks reported as Avg.
+DitHub demonstrates strong performance on both average and individual tasks, consistently outperforming its competitor by a notable margin.
+Given the limited class overlap within ODinW-13, the primary component susceptible to catastrophic forgetting in this analysis is the shared matrix $B$.
+Notably, for some tasks, this metric is even positive for our approach.
+We attribute this behavior to the initialization of the $B$ component.
+Specifically, the metric is higher in the first learned task (Pa)  [^4] The order in which tasks are presented to the model is random and dependent on the selected seed..
+While this might seem counterintuitive, the Pa dataset comprises very few instances, which means the $B$ matrix cannot undergo robust initialization.
+Consequently, subsequent tasks introduce greater data diversity, leading to improved performance compared to the initial one limited by scarce training data.
+B.2 Forgetting on ODinW-O
+Table D: Forgetting metrics in terms of mAP for the ODinW-O dataset. Bold indicates best results.
+$5$ $6$ $1$ $2$ $4$ $3$
+Method Avg Ae Hw Pv Sd Th Ve
+ZiRa -$\texttt{-}3.84$$\boldsymbol{\texttt{-}0.08}$ $0.00$$\texttt{-}9.67$$\texttt{-}10.66$$\texttt{-}0.98$$\texttt{-}1.62$
+DitHub -$\boldsymbol{\texttt{-}2.68}$$\texttt{-}4.75$ $0.00$$\boldsymbol{\texttt{-}2.74}$$\boldsymbol{\texttt{-}9.51}$ $\boldsymbol{~{}~{}1.74}$$\boldsymbol{\texttt{-}0.79}$
+Tab.
+D reports the forgetting metric results for the ODinW-O benchmark.
+Unlike the previous analysis, this assessment allows us to evaluate the forgetting of DitHub’s class-specific $A$ components for all classes within the dataset.
+Indeed, ODinW-O is specifically designed to test Incremental Learning capabilities, featuring complete class overlap across tasks.
+This overlap leads to catastrophic forgetting of the $A_{c}$ expert component for the same class $c$ across different domains.
+It’s worth noting that even in this setting, which emphasizes the $A$ component, the forgetting of $B$ remains a concern.
+Even in this challenging context, DitHub surpasses its competitor by an appreciable margin, demonstrating difficulty only with the Ae task and highlighting the non-trivial nature of merging and preserving knowledge from the aerial domain.
+Appendix C Additional Benchmark Details
 This section provides additional details regarding the datasets utilized in our evaluation framework.
@@
 By leveraging these datasets, we establish a rigorous evaluation protocol to comprehensively assess model robustness in challenging real-world scenarios.
-B.1 ODinW-13 and MS COCO
+C.1 ODinW-13 and MS COCO
 Following [4], we assess our model specialization using the Object Detection in the Wild (ODinW-13) [20] benchmark and the MS COCO [22] dataset to verify whether the zero-shot capabilities are preserved after the sequential fine-tuning on ODinW-13.
@@
 This comprehensive collection covers rare object categories, diverse contextual conditions, and challenging domains.
-Consequently, the ODinW-13 benchmark offers an ideal scenario to evaluate the capability of DitHub to sequentially adapt to heterogeneous task distributions while effectively mitigating catastrophic forgetting on previously learned tasks.
+Consequently, the ODinW-13 benchmark offers an ideal scenario to evaluate the capability of DitHub to sequentially adapt to heterogeneous task distributions.
 Following the convention established by [4], we use the term ZCOCO to denote the zero-shot evaluation conducted on the MS COCO validation set comprising $5000$ samples and consisting of annotations for $80$ common object categories.
 The ZCOCO evaluation specifically measures the robustness and generalization capabilities of DitHub in zero-shot scenarios, after training on each successive task from ODinW-13.
-B.2 ODinW-O
+C.2 ODinW-O
+Person Car Truck Dog Boat
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure B: Visualization of domain shifts for shared categories across different tasks within the ODinW-O benchmark.
 We introduce a new benchmark, ODinW-O (Overlapped), designed to further evaluate the class-incremental and specialization capabilities of DitHub.
-Derived from ODinW-35, ODinW-O retains only classes shared across at least one task (overlapping classes), while removing tasks with no shared classes.
-This filtering is necessary because classes that do not reappear in subsequent tasks present an ideal case in the DitHub framework.
-Specifically, their corresponding non-recurring class modules remain frozen, requiring no further adaptation.
-In contrast, ODinW-O eliminates this scenario by ensuring that all retained classes appear in multiple tasks.
-By enforcing strictly overlapping classes, ODinW-O provides a challenging benchmark for evaluating incremental learning, specifically tailored to assess the robustness of our method.
+Derived from ODinW-35 [19], ODinW-O retains only classes shared across at least one task (i.e., overlapping classes), while removing tasks with no shared classes.
+This filtering is crucial because retaining classes that do not reappear in subsequent tasks would cause the benchmark to resemble the original one.
+Specifically, the corresponding non-recurring class modules would remain frozen, requiring no further adaptation.
+In contrast, ODinW-O eliminates this scenario by ensuring all retained classes appear in multiple tasks.
+By enforcing strictly overlapping classes, ODinW-O provides a challenging benchmark for evaluating Incremental Learning.
 Additionally, to rigorously evaluate the model robustness to domain shifts, we designed ODinW-O by explicitly selecting datasets that feature markedly different contexts (Fig.
 B), enabling clear insights into how effectively the model adapts to variations in visual appearance, environmental conditions, and imaging domains.
-We select the following datasets, already presented in Sec.
-B.1, and extracted only the overlapping classes:
-  • Aerial Maritime Drone: boat and car,
-  • Pascal VOC: boat, car, dog, and person
-  • Thermal Dogs and People: dog and person,
-  • Vehicles: car and truck,
+We select the following datasets, already mentioned in Sec.
+C.1, and extracted only the overlapping classes:
+  • Aerial Maritime Drone: boat and car;
+  • Pascal VOC: boat, car, dog, and person;
+  • Thermal Dogs and People: dog and person;
+  • Vehicles: car and truck.
 Additionally, we introduce two further datasets from ODinW-35 to emphasize domain variation:
-  • Hard Hat Workers: person
-  • Self Driving Cars: car and truck
+  • Hard Hat Workers: person;
+  • Self Driving Cars: car and truck.
 Despite ODinW-O enforcing class overlap and introducing domain shifts, our model consistently achieves state-of-the-art performance by effectively refining class knowledge while adapting across diverse contexts.
-Person Car
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure B: Visualization of domain shifts for shared categories across different tasks within the ODinW-O benchmark.
-Appendix C Implementation Details
-All experiments for DitHub were conducted on a single RTX A5000 GPU with 24GB of VRAM, requiring approximately 8 hours per training run with a batch size of 2.
-During training, our implementation in PyTorch efficiently utilizes around 10 GB of VRAM.
-We employed GroundingDINO-T (Swin-T) as an open-vocabulary object detector, pretrained on O365 [39], GoldG [17], and Cap4M [17] datasets.
-We performed training with random shuffling across three different seeds to ensure robustness.
-For LoRA fine-tuning, we perform low-rank updates to the attention layers with a rank of 16, applying a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-2}$.
-This approach enables efficient adaptation while significantly reducing computational overhead and memory storage for each class-specialized module.
-Optimization is performed using AdamW.
-We rely solely on the contrastive classification loss and the localization loss of Grounding DINO, without introducing any additional loss functions.
-The full Grounding DINO model, after integrating 44 classes during training on ODinW-13, consists of approximately 173 million parameters, with 19.6 million LoRA parameters actively updated during training.
-After fine-tuning on ODinW-13, only 75.65 MB of LoRA parameters are retained for storage.
 Appendix D Additional Results
+Figure C: Per-class mAP degradation after removing class modules. Darker cells indicate greater performance drops.
+D.1 Standard Deviations
+We report the standard deviations of the reproduced methods from Tab. 1 and Tab. 2 in Tab. E and F, respectively, computed over three runs with different random seeds.
+D.2 Unlearning Additionals
+As previously discussed in Sec.
+5.2, our modular framework facilitates the precise removal of specific knowledge from Grounding DINO through the subtraction of corresponding experts.
+Fig.
+C visually represents this process: each column depicts the mean Average Precision (mAP) of a target class after unlearning different modules (reported on the y-axis), with colors normalized column-wise to emphasize relative performance decrements.
+The prominent darker diagonal pattern signifies that removing a class-specific module results in the most substantial performance reduction for its associated class.
+This observation substantiates that our modules effectively encapsulate specialized knowledge, as their unlearning selectively impacts their respective classes.
+Furthermore, this matrix representation elucidates correlations between certain classes during the unlearning procedure.
+For instance, the darker shading in the “bicycle” column indicates that the "person" class exhibits the second most significant mAP drop.
+This correlation arises from the frequent co-occurrence of “person” instances riding bicycles, naturally linking these classes within the comprehensive unlearning pipeline.
+Table E: Standard deviations of the reproduced methods in Tab. 1.
+Shots Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
+1 ZiRa ±$\pm 0.90$$\pm 2.05$$\pm 1.49$$\pm 0.17$$\pm 2.16$$\pm 0.98$$\pm 0.79$$\pm 1.63$$\pm 1.33$$\pm 7.91$$\pm 3.29$$\pm 12.17$$\pm 4.45$$\pm 3.07$$\pm 3.25$
+DitHub ±$\pm 0.60$$\pm 0.81$$\pm 0.95$$\pm 0.57$$\pm 2.66$$\pm 1.58$$\pm 3.72$$\pm 0.00$$\pm 1.53$$\pm 1.10$$\pm 2.90$$\pm 7.03$$\pm 3.68$$\pm 4.66$$\pm 2.46$
+5 ZiRa ±$\pm 0.23$$\pm 0.43$$\pm 3.76$$\pm 2.15$$\pm 1.40$$\pm 1.08$$\pm 3.54$$\pm 3.74$$\pm 1.54$$\pm 4.26$$\pm 3.73$$\pm 4.13$$\pm 2.03$$\pm 2.02$$\pm 4.16$
+DitHub ±$\pm 0.42$$\pm 1.27$$\pm 2.52$$\pm 0.70$$\pm 2.42$$\pm 6.74$$\pm 0.76$$\pm 3.27$$\pm 0.17$$\pm 5.49$$\pm 4.79$$\pm 2.22$$\pm 4.29$$\pm 1.81$$\pm 1.54$
+10 ZiRa ±$\pm 0.28$$\pm 0.66$$\pm 0.44$$\pm 2.64$$\pm 2.77$$\pm 0.52$$\pm 6.43$$\pm 0.97$$\pm 1.04$$\pm 2.17$$\pm 2.04$$\pm 3.90$$\pm 1.21$$\pm 3.71$$\pm 0.87$
+DitHub ±$\pm 0.20$$\pm 0.54$$\pm 2.26$$\pm 0.72$$\pm 4.23$$\pm 3.53$$\pm 1.35$$\pm 3.19$$\pm 0.30$$\pm 2.74$$\pm 0.34$$\pm 2.99$$\pm 1.01$$\pm 1.15$$\pm 1.27$
+Full ZiRA ±$\pm 0.44$$\pm 0.68$$\pm 2.73$$\pm 1.17$$\pm 1.39$$\pm 4.77$$\pm 7.26$$\pm 2.19$$\pm 1.33$$\pm 0.33$$\pm 4.14$$\pm 3.57$$\pm 2.44$$\pm 2.82$$\pm 3.13$
+DitHub ±$\pm 0.16$$\pm 0.44$$\pm 2.39$$\pm 1.00$$\pm 2.83$$\pm 0.48$$\pm 1.34$$\pm 0.60$$\pm 0.38$$\pm 1.68$$\pm 0.76$$\pm 1.57$$\pm 2.87$$\pm 0.69$$\pm 1.01$
+Table F: Standard deviations of the reproduced methods in Tab. 2.
 Method ZCOCO Avg Ae Hw Pv Sd Th Ve
 ZiRa ±$\pm 0.25$$\pm 0.27$$\pm 1.76$$\pm 1.14$$\pm 1.35$$\pm 0.32$$\pm 1.75$$\pm 1.76$
-Ours ±$\pm 0.24$$\pm 0.66$$\pm 2.39$$\pm 1.88$$\pm 0.76$$\pm 1.31$$\pm 0.46$$\pm 0.84$
-Table C: Standard deviations of the reproduced methods in Table 2.
-D.1 Standard Deviations
-Shots Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
-1 ZiRa $\pm$ 0.90 $\pm$ 2.05 $\pm$ 1.49 $\pm$ 0.17 $\pm$ 2.16 $\pm$ 0.98 $\pm$ 0.79 $\pm$ 1.63 $\pm$ 1.33 $\pm$ 7.91 $\pm$ 3.29 $\pm$ 12.17 $\pm$ 4.45 $\pm$ 3.07 $\pm$ 3.25
-DitHub $\pm$ 0.60 $\pm$ 0.81 $\pm$ 0.95 $\pm$ 0.57 $\pm$ 2.66 $\pm$ 1.58 $\pm$ 3.72 $\pm$ 0.00 $\pm$ 1.53 $\pm$ 1.10 $\pm$ 2.90 $\pm$ 7.03 $\pm$ 3.68 $\pm$ 4.66 $\pm$ 2.46
-5 ZiRa $\pm$ 0.23 $\pm$ 0.43 $\pm$ 3.76 $\pm$ 2.15 $\pm$ 1.40 $\pm$ 1.08 $\pm$ 3.54 $\pm$ 3.74 $\pm$ 1.54 $\pm$ 4.26 $\pm$ 3.73 $\pm$ 4.13 $\pm$ 2.03 $\pm$ 2.02 $\pm$ 4.16
-DitHub $\pm$ 0.42 $\pm$ 1.27 $\pm$ 2.52 $\pm$ 0.70 $\pm$ 2.42 $\pm$ 6.74 $\pm$ 0.76 $\pm$ 3.27 $\pm$ 0.17 $\pm$ 5.49 $\pm$ 4.79 $\pm$ 2.22 $\pm$ 4.29 $\pm$ 1.81 $\pm$ 1.54
-10 ZiRa $\pm$ 0.28 $\pm$ 0.66 $\pm$ 0.44 $\pm$ 2.64 $\pm$ 2.77 $\pm$ 0.52 $\pm$ 6.43 $\pm$ 0.97 $\pm$ 1.04 $\pm$ 2.17 $\pm$ 2.04 $\pm$ 3.90 $\pm$ 1.21 $\pm$ 3.71 $\pm$ 0.87
-DitHub $\pm$ 0.20 $\pm$ 0.54 $\pm$ 2.26 $\pm$ 0.72 $\pm$ 4.23 $\pm$ 3.53 $\pm$ 1.35 $\pm$ 3.19 $\pm$ 0.30 $\pm$ 2.74 $\pm$ 0.34 $\pm$ 2.99 $\pm$ 1.01 $\pm$ 1.15 $\pm$ 1.27
-Full ZiRA $\pm$ 0.44 $\pm$ 0.68 $\pm$ 2.73 $\pm$ 1.17 $\pm$ 1.39 $\pm$ 4.77 $\pm$ 7.26 $\pm$ 2.19 $\pm$ 1.33 $\pm$ 0.33 $\pm$ 4.14 $\pm$ 3.57 $\pm$ 2.44 $\pm$ 2.82 $\pm$ 3.13
-DitHub $\pm$ 0.16 $\pm$ 0.44 $\pm$ 2.39 $\pm$ 1.00 $\pm$ 2.83 $\pm$ 0.48 $\pm$ 1.34 $\pm$ 0.60 $\pm$ 0.38 $\pm$ 1.68 $\pm$ 0.76 $\pm$ 1.57 $\pm$ 2.87 $\pm$ 0.69 $\pm$ 1.01
-Table D: Standard deviations of the reproduced methods in Table 1.
-We report the standard deviations of the reproduced methods from Tab. 1 and Tab. 2 in Tab. D and C, respectively, computed over three runs with different random seeds.
-D.2 Unlearning Additionals
-Figure C: Matrix visualization of mAP degradation per class after removing class modules. Colors are normalized column-wise for readability; darker cells indicate greater performance drops.
-As discussed in Sec.
-5.2, our modular approach enables the precise removal of specific knowledge from Grounding DINO by subtracting corresponding experts.
-Figure C illustrates this process: each column represents a target class’s mAP after unlearning different modules, with colors normalized column-wise to highlight relative drops.
-The darker diagonal pattern indicates that removing a class module leads to the most significant performance drop for its corresponding class.
-This confirms that our modules effectively encapsulate specialized knowledge, as unlearning them selectively impacts their respective classes.
-Appendix E Qualitatives
+DitHub ±$\pm 0.24$$\pm 0.66$$\pm 2.39$$\pm 1.88$$\pm 0.76$$\pm 1.31$$\pm 0.46$$\pm 0.84$
+Appendix E Implementation Details
+All experiments for DitHub were conducted on a single RTX A5000 GPU with 24 GB of VRAM.
+Each training run, with a batch size of $2$, required approximately $8$ hours.
+Our PyTorch implementation efficiently utilized around 10 GB of VRAM during training.
+We employed Grounding DINO Tiny as an Open-Vocabulary object detector, which was pre-trained on the O365 [44], GoldG [16], and Cap4M [20] datasets.
+For LoRA fine-tuning, we applied low-rank updates to the encoder layers with a rank of $16$.
+The optimization was performed using AdamW [26], with a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-2}$.
+We exclusively relied on Grounding DINO’s contrastive classification loss and localization loss, without introducing any additional loss functions.
+The final model comprised approximately $173$ million parameters, with only $19.6$ million LoRA parameters actively updated during training on ODinW-13.
+After fine-tuning on ODinW-13, only 75.65 MB of LoRA parameters were retained for storage.
+Appendix F Limitations and Broader Impact
+A limitation of the current work is that the merging strategies employed by DitHub for both class-specific experts ($A$ matrices) and the shared $B$ matrix could be further optimized.
+While the current solutions demonstrated effectiveness, their reliance on empirically derived coefficients highlights a potential for refinement towards strategies grounded in more robust theoretical principles.
+Moreover, when merging modules pertaining to the same class but originating from different tasks, incorporating explicit domain knowledge becomes crucial.
+We hypothesize that a mechanism explicitly accounting for domain information (e.g., distinguishing between RGB and thermal images) could not only enhance DitHub’s performance but also facilitate the development of novel expert module merging techniques.
+Despite these acknowledged limitations, DitHub achieves state-of-the-art performance, underscoring the efficacy of its expert module management approach.
+We propose that our method, and more broadly Modular Deep Learning, can serve as a cornerstone for fostering more accessible Artificial Intelligence.
+The development of methods for training and managing libraries of expert modules could enable the granular selection and deployment of knowledge onto large, otherwise computationally intractable, pre-trained architectures.
+Specifically, DitHub facilitates the management of modules that enable large Open-Vocabulary object detectors to recognize specific and potentially rare classes.
+This specificity also contributes to enhanced model controllability, potentially enabling targeted unlearning capabilities for bias mitigation and thereby advancing towards fairer deep learning systems.
+Appendix G Qualitative Results
 In Fig.
@@
 Our method excels at detecting challenging objects, including small ones that other models miss.
-Additionally, DitHub avoids a false positive (ZiRA) and a false negative (Grounding DINO), demonstrating more reliable predictions.
+Additionally, DitHub avoids false positives (ZiRA) and false negatives (Grounding DINO), demonstrating more reliable predictions.
 It also correctly identifies both cars (last row) despite occlusions and distractors.

### v3

@@
 At inference, it enables the selective activation of specialized modules for rare or complex classes, ensuring robust adaptation while maintaining the backbone’s zero-shot and Open-Vocabulary capabilities.
-A high-level overview of our approach is depicted in Footnote 2.
+A high-level overview of our approach is depicted in Footnote˜2.
 Following [4], we evaluate DitHub on the ODinW-13 [20] benchmark, where it outperforms the state-of-the-art by a substantial $\boldsymbol{\texttt{+}4.21}$ mAP in the Incremental Vision-Language Object Detection setting.
@@
 Formally, for a given task $t$, we seek an adaptation $\Delta$ that minimizes the following objective:
-$\min_{\Delta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\;\ell(f(x;(\Theta,\Delta)),\,y)+\ell_{\text{zero-shot}},$ (1)
+$\min_{\Delta}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\;\ell(f(x;\Theta+\Delta),\,y)+\ell_{\text{zero-shot}},$ (1)
 where $\ell$ is the Object Detection loss function.
-The term $(\Theta,\Delta)$ represents the set of adapted model parameters.
+The term $\Theta+\Delta$ represents the set of adapted model parameters.
 Additionally, $\ell_{\text{zero-shot}}$ quantifies the performance difference in the zero-shot setting before and after adaptation.
@@
 LoRA.
-Given a layer with weight $W\in\mathbb{R}^{d\times k}$, Low-Rank Adaptation [13] (LoRA) uses low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ – with $r\ll\min(d,k)$ – such that the adapted weight $W^{\prime}=W+\Delta W=W+BA$.
-The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to full fine-tuning.
+Low-Rank Adaptation [13] is a technique to significantly reduce the number of trainable parameters.
+Given a generic layer with weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA introduces low-rank matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$ — with $r\ll\min(d,k)$ — such that the adapted weight $W^{\prime}$ is:
+$W^{\prime}=W+\Delta W,\quad\text{where}\quad\Delta W=BA.$ (2)
+The use of these low-rank matrices significantly reduces memory usage and computational overhead, making LoRA an efficient alternative to fine-tuning the entire model.
 4 DitHub
@@
 Inspired by this, we propose DitHub — a framework that operates akin to a Version Control System, enabling the creation and expansion of a library of efficient adaptation modules for Vision-Language object detectors.
-DitHub develops and maintains class-specific expert modules using a simple yet effective strategy that requires no architectural modifications or additional loss terms (Sec.
-4.1).
+DitHub develops and maintains class-specific expert modules using a simple yet effective strategy that requires no architectural modifications or additional loss terms (Section˜4.1).
 Following [4], we consider the pre-trained Grounding DINO as the base architecture for Open-Vocabulary detection and use LoRA for fine-tuning.
-We propose an efficient pipeline that handles LoRA’s low-rank matrices with two objectives: matrix $A$ encodes class-specific knowledge, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency (Sec.
-4.2).
-Finally, when fine-grained adaptation is needed (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Sec.
-4.3.
-The pseudocode of DitHub is provided in Alg.
-1.
+We propose an efficient pipeline that handles LoRA’s low-rank matrices with two objectives: matrix $A$ encodes class-specific knowledge, while matrix $B$ captures general knowledge and is shared among tasks to maintain overall efficiency (Section˜4.2).
+Finally, when fine-grained adaptation is needed (e.g., detecting a rare object within a scene), DitHub enables the dynamic extraction of the corresponding expert modules, ensuring precise and efficient inference, as outlined in Section˜4.3.
+The pseudocode of DitHub is provided in Algorithm˜1.
 4.1 Training Class-Specific Modular Detectors
@@
 However, preliminary experiments show this approach is ineffective for rare categories, as random selection limits their exposure to sufficient training data.
-As better highlighted in Alg.
-1, we overcome this issue by introducing an earlier warmup phase at the beginning of each task.
+As better highlighted in Algorithm˜1, we overcome this issue by introducing an earlier warmup phase at the beginning of each task.
 This phase is meant to provide a robust, class-agnostic initialization, ensuring that the subsequent specialization stage is more effective.
 Warmup.
-At the onset of the new task, a shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Fig.
-2, yellow arrow) to capture class-agnostic yet task-specific knowledge.
-This phase, run for a fixed number of epochs, provides a common initialization for the $A$ modules before specialization.
+At the onset of the new task, a shared warmup matrix $A_{\text{wu}}$ is instantiated and optimized (see Figure˜2, yellow arrow) to capture class-agnostic yet task-specific knowledge.
+This phase, which runs for a fixed number of epochs, provides a common initialization for the $A$ modules before their specialization.
 By doing so, the specialized modules start from a common base point, a factor that has been shown to greatly enhance model compositionality [33, 39].
-Often associated with improved linear mode connectivity between models [9], these properties are leveraged by our approach to combine class-specific modules based on input textual queries, as detailed in Sec.
-4.3.
+Often associated with improved linear mode connectivity between models [9], these properties are leveraged by our approach to combine class-specific modules based on input textual queries, as detailed in Section˜4.3.
 Moreover, considering rare and underrepresented classes with a few training examples, the warmup phase seeks to establish an already robust and effective initialization.
@@
 After the warmup phase, we proceed to the core of DitHub: the specialization phase.
-To train class-specific modules, DitHub generates $|\mathcal{C}_{t}|$ distinct branches (branch in Fig.
-2), each dedicated to a single class.
-These branches undergo independent optimization, represented by a blue arrow in Fig.
-2, to capture class-specific capabilities.
-In doing so, for each training image, we employ the stochastic strategy introduced in Sec.
-4.1 to randomly select a single expert for training.
+To train class-specific modules, DitHub generates $|\mathcal{C}_{t}|$ distinct branches (branch in Figure˜2), each dedicated to a single class.
+These branches undergo independent optimization, represented by a blue arrow in Figure˜2, to capture class-specific capabilities.
+In doing so, for each training image, we employ the stochastic strategy introduced in Section˜4.1 to select a single expert for training.
 Notably, the modular design of our approach effectively handles potential class overlaps across consecutive tasks.
@@
    if $A_{c}\in\mathcal{A}$ then $\triangleright$  Fetch
-              $A_{c}^{\text{cur}}\leftarrow$ use Eq. 2 $\triangleright$  Merge
+     $A_{c}^{\text{cur}}\leftarrow$ use Equation˜3 $\triangleright$  Merge
    else
@@
    $\mathcal{A}\leftarrow\mathcal{A}\cup A_{c}$
-     $B\leftarrow$ use Eq. 3
+  $B\leftarrow$ use Equation˜4
 Figure 3: Memory requirements as the number of classes grows.
-To favor selective updates, DitHub performs a fetch operation (Fig.
-2), which simply retrieves the class-specific module from previous tasks if the class has already been seen — followed by the merge (Fig.
-2) .
+To favor selective updates, DitHub performs a fetch operation (Figure˜2), which simply retrieves the class-specific module from previous tasks if the class has already been seen — followed by the merge (Figure˜2) .
 For instance, considering the class “dog”, the previously stored expert module $A_{\text{dog}}^{\text{old}}$ is fetched and merged with the current task’s warmup matrix $A_{\text{wu}}$ following a weighted strategy:
@@
 Here, $\lambda_{A}$ is set to a low value, prioritizing consolidated knowledge over new insights from the warmup phase.
-Details on the impact of $\lambda_{A}$ are in Sec.
-5.2.
+Details on the impact of $\lambda_{A}$ are in Section˜5.2.
 Crucially, our merging process creates an on-the-fly initialization that blends historical class-specific knowledge with new domain information, resulting in $A_{\text{dog}}^{\text{cur}}$, which serves as the starting point for the specialization phase.
@@
 To keep our library memory-efficient, we maintain a single shared $B$ matrix instead of a distinct $B$ for each class.
-As shown in Fig.
-3, using class-specific $(A,B)$ pairs results in a linear increase in memory consumption as the number of classes grows.
-Sharing the $B$ matrix reduces this requirement by half  [^3] In Sec.
-A.2, we present DitHub’s results when $B$ is not shared..
+As shown in Figure˜3, using class-specific $(A,B)$ pairs results in a linear increase in memory consumption as the number of classes grows.
+Sharing the $B$ matrix reduces this requirement by half  [^3] In Section A.2, we present DitHub’s results when $B$ is not shared..
 Additionally, considering DitHub “$B$ shared” (line [Uncaptioned image]), the memory footprint of our approach is consistently lower than the storage needed for the Grounding DINO encoder alone (line [Uncaptioned image]).
 This makes DitHub’s memory demands manageable for practical applications.
-Moreover, memory efficiency is further optimized by fine-tuning a targeted portion of the base architecture (Sec.
-A.1).
+Moreover, memory efficiency is optimized by fine-tuning a targeted portion of the base architecture (Section˜A.1).
 Due to the shared design, the matrix $B$ is more susceptible to interference and catastrophic forgetting [31].
@@
 $B_{t}=(1-\lambda_{B})B_{t-1}+\lambda_{B}B^{\text{opt}}.$ (4)
-Unlike Eq.
-2, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state.
-This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, further detailed in Sec.
-5.2.
+Unlike Equation˜3, the merging coefficient $\lambda_{B}$ is assigned a higher value, giving greater weight to the latest state.
+This approach results in a dual strategy for $\lambda_{A}$ and $\lambda_{B}$, further detailed in Section˜5.2.
 4.3 Inference with DitHub
@@
 To better assess incremental detection and module reuse, we introduce ODinW-O (Overlapped), a subset of ODinW-35 that is designed to reflect realistic scenarios where classes can reoccur across tasks over time.
-Further details are provided in App.
-C.
+Further details are provided in Appendix˜C.
 For comparisons, we employ mean Average Precision (mAP) and compare DitHub against six state-of-the-art methods.
@@
 Implementation details.
-We focus on Grounding DINO and adapt only its encoder network, as further validated in Sec.
-A.1.
+We focus on Grounding DINO and adapt only its encoder network, as further validated in Section˜A.1.
 The merging coefficient $\lambda$ for the $A$ matrix is set to $0.3$ for ODinW-13 and $0.1$ for ODinW-O; for the matrix $B$, we set $\lambda=0.7$.
-The choice of $\lambda$ and its underlying rationale are further discussed in Sec.
-5.2.
-During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — see App.
-E for further hyperparameter details.
+The choice of $\lambda$ and its underlying rationale are further discussed in Section˜5.2.
+During training, we allocate an equal number of epochs to the warmup and the subsequent specialization phases — see Appendix˜E for further details.
 Performance on ODinW-13.
-Tab.
-1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO).
-We evaluate performance in both the full-data setting and few-shot scenarios with $1$, $5$, and $10$ shots per class.
+Table˜1 presents the IVLOD results, including the average across the $13$ tasks (Avg) and zero-shot preservation (ZCOCO).
+We evaluate performance in both the full-data setting and few-shot scenarios with $1$, 55, and $10$ shots per class.
 All baseline results, except those of Grounding DINO, ZiRa, and our method, are taken directly from [4].
-Each number represents the average of three runs with different random seeds, with standard deviations provided in Sec.
-D.1.
+Each number represents the average of three runs with different random seeds, with standard deviations provided in Section˜D.1.
 Table 1: Comparison of mAP values across ODinW-13 tasks. Avg represents the average across the $13$ tasks, while zero-shot performance is reported in the ZCOCO column. Bold indicates best results.
@@
 DitHub $\boldsymbol{47.01}$ $\boldsymbol{62.19}$ $\boldsymbol{34.62}$ $\boldsymbol{50.65}$ $70.46$ $68.56$ $49.28$ $\boldsymbol{65.57}$ $\boldsymbol{69.58}$ $71.10$ $\boldsymbol{56.65}$ $\boldsymbol{70.88}$ $\boldsymbol{52.82}$ $\boldsymbol{79.30}$ $\boldsymbol{68.18}$
+Figure 4: Normalized Average mAP as tasks progress.
+To assess performance degradation, we normalize the mAP for each task by dividing the current value by the original mAP recorded at the end of that task, i.e., if the curve $\approx 0.90$, the model retains $90\%$ of its original performance.
 In the full-shot setting, DitHub outperforms ZiRa by a substantial $\boldsymbol{\texttt{+}4.21}$ mAP and achieves the best results in nine out of thirteen tasks.
@@
 The ODinW-O scenario emphasizes that frequently encountered classes may reappear in subsequent tasks or at later times, such as the class “person” in common industrial and smart-city scenarios.
-As detailed in Sec.
-C.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks.
+As detailed in Section˜C.2, ODinW-O guarantees that each class is revisited at least once in subsequent tasks.
 In this benchmark, we primarily compare our approach with ZiRa, as it is the most competitive method in the standard IVLOD setting, excluding ours.
@@
 DitHub $\boldsymbol{46.51}$ $\boldsymbol{62.38}$ $\boldsymbol{53.35}$ $\boldsymbol{71.07}$ $\boldsymbol{71.01}$ $41.75$ $\boldsymbol{80.21}$ $\boldsymbol{56.90}$
-The results in Tab.
-2 indicate that our approach secures a substantial improvement of $\boldsymbol{\texttt{+}4.75}$ in mAP and a $\boldsymbol{\texttt{+}2.08}$ mAP gain on ZCOCO over ZiRa.
+The results in Table˜2 indicate that our approach secures a substantial improvement of $\boldsymbol{\texttt{+}4.75}$ in mAP and a $\boldsymbol{\texttt{+}2.08}$ mAP gain on ZCOCO over ZiRa.
 We attribute the success of our method to its class-oriented modular design, which enables selective updates to recurring concepts.
 This design prevents unintentional overwriting of knowledge across classes, tasks, and domains.
-Figure 4: Normalized Average mAP as tasks progress.
-To assess performance degradation, we normalize the mAP for each task by dividing the current value by the original mAP recorded at the end of that task, i.e., if the curve $\approx 0.90$, the model retains $90\%$ of its original performance.
 Performance analysis as tasks progress.
-In Fig.
-4, we present an incremental analysis comparing the performance of DitHub and its strongest competitor, ZiRa, after each task.
-The evaluation, conducted on both ODinW-13 (left) and ODinW-O (right), allows us to assess the retention of both zero-shot and in-domain capabilities throughout the task sequence, providing a direct measure of their resistance to forgetting (further discussed in App.
-B).
+In Figure˜4, we present an incremental analysis comparing the performance of DitHub and its strongest competitor, ZiRa, after each task.
+The evaluation, conducted on both ODinW-13 (left) and ODinW-O (right), allows us to assess the retention of both zero-shot and in-domain capabilities throughout the task sequence, providing a direct measure of their resistance to forgetting (further discussed in Appendix˜B).
 On ZCOCO, DitHub exhibits stable performance across tasks, with a negligible loss in zero-shot capabilities — unlike ZiRa, which shows a more pronounced decline.
@@
 To this aim, we introduce an ablative approach, Experts-not-Experts (EnE), which removes class specialization.
-EnE is identical to DitHub, except during training, where it omits the stochastic strategy from Sec.
-4.1.
+EnE is identical to DitHub, except during training, where it omits the stochastic strategy from Section˜4.1.
 Instead, given an image from the current task, EnE performs the training step on one module chosen at random among all task classes.
@@
 Consequently, they become generalist learners rather than class-specific experts.
-In Tab.
-3, we compare DitHub with its non-specialized variant, i.e., EnE.
+In Table˜3, we compare DitHub with its non-specialized variant, i.e., EnE.
 Notice that, at inference, EnE acts as an ensemble of class-agnostic modules.
@@
 DitHub $47.01$ $62.19$ $34.62$ $50.65$ $70.46$ $68.56$ $49.28$ $65.57$ $69.58$ $71.10$ $56.65$ $70.88$ $52.82$ $79.30$ $68.18$
-Figure 5: Results on rare classes of ODinW-13.
-Figure 6: Showing unlearning skills of DitHub.
+Figure 5: Performance comparison of Grounding DINO, EnE, and DitHub on rare object categories from the ODinW-13 dataset.
+Figure 6: Sensitivity analysis on $\lambda_{A}$ and $\lambda_{B}$. Performance measured as harmonic mean of Avg and ZCOCO mAP.
 Evaluation on rare classes.
-As discussed in Sec.
-4.3, DitHub offers fine-grained control at inference, enabling users to tailor the system to their specific use case — including the option to use a single module specialized for a target class.
+As discussed in Section˜4.3, DitHub offers fine-grained control at inference, enabling users to tailor the system to their use case — including the option to use a single module specialized for a target class.
 To assess this, we conduct a targeted evaluation on rare classes from ODinW-13, activating only the corresponding target class module.
-Results in Fig.
-6 show that DitHub achieves the best results on rare classes, highlighting the benefits of class specialization.
-Specialized modules allows training-free unlearning.
-DitHub’s advantages go beyond additive Incremental Learning.
-Expert modules can also be used to remove specific knowledge by simply subtracting the corresponding weights from the base model.
-This operation, akin to zero-training unlearning, has been studied in image and text categorization [15, 33]; to the best of our knowledge, we are the first to explore it within the context of Object Detection.
-For example, this enables the model to unlearn the detection of sensitive object categories – such as faces, license plates, or medical equipment – when required to comply with privacy regulations or data retention policies.
-Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class $c$ module from the pre-trained model: $W^{\prime}=W-\alpha BA_{c}$, whit $\alpha=0.3$.
-Fig.
-6 illustrates this qualitatively: on the left, the zero-shot model correctly detects common classes, while on the right, subtracting a class module reduces its mAP, effectively unlearning it and leading to false negatives.
-A quantitative analysis is provided in Sec.
-D.2.
+Results in Figure˜6 show that DitHub achieves the best results on rare classes, highlighting the benefits of class specialization.
 Sensitivity Analysis of $\boldsymbol{\lambda}$.
@@
 To this end, we define an aggregate metric as the harmonic mean of DitHub’s full-shot performance, balancing specialization (Avg) and zero-shot generalization (ZCOCO).
-Fig.
-8 shows performance trends for ODinW-13 (top) and ODinW-O (bottom) as $\lambda_{A}$ and $\lambda_{B}$ in $0.2$ increments.
-Considering $\lambda_{A}$, its role is significant while merging the expert $A$ matrices (see Eq.
-2).
-From Fig.
-8, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective.
+Figure˜6 shows performance trends for ODinW-13 (top) and ODinW-O (bottom) as $\lambda_{A}$ and $\lambda_{B}$ in $0.2$ increments.
+Considering $\lambda_{A}$, its role is significant while merging the expert $A$ matrices (see Equation˜3).
+From Figure˜6, lower values of $\lambda_{A}$ lead to better results, indicating that prioritizing past expert knowledge over the warmup module is more effective.
 Interestingly, while $\lambda_{A}=0.3$ performs best for ODinW-13, ODinW-O exhibits a monotonic decline as $\lambda_{A}$ increases.
@@
 The harmonic mean increases with $\lambda_{B}$, peaking at $0.7$ for both datasets.
-This suggests that prioritizing the knowledge in the $B^{\text{opt}}$ matrix from the current task, as shown in Eq.
-3, is crucial.
+This suggests that prioritizing the knowledge in the $B^{\text{opt}}$ matrix from the current task, as shown in Equation˜4, is crucial.
 Our intuition, supported by [30], is that the latest $B^{\text{opt}}$ matrix integrates knowledge from both earlier and recent tasks, making it more valuable than $B_{t-1}$, which is trained solely on earlier tasks.
-Figure 7: Sensitivity analysis on $\lambda_{A}$ and $\lambda_{B}$. Performance measured as harmonic mean of Avg and ZCOCO mAP.
-Figure 8: Ablative study on the DitHub components. Results are expressed as average performance across the tasks of ODinW-13.
+Method Memory (MB) Avg (mAP)
+DitHub $r=1$ $\sim 9$ $57.04$
+ZiRa $\sim 18$ $57.98$
+DitHub $r=2$ $\sim 18$ $60.26$
+DitHub $r=4$ $\sim 37$ $60.53$
+DitHub $r=8$ $\sim 74$ $61.93$
+DitHub $r=16$ $\sim 147$ $62.19$
+G-Dino (encoder only) $\sim 659$ $46.80$
+Figure 7: (left) Ablation study on DitHub components, showing average performance on ODinW-13 alongside the ZCOCO metric.
+(right) Memory usage (MB) versus performance (mAP) for DitHub at varying ranks, compared against ZiRa, with performance averaged over the ODinW-13 tasks.
 On Warmup and Shared $\boldsymbol{B}$.
-To evaluate the impact of DitHub’s components, we progressively integrate them and present results in Fig.
-8.
+To evaluate the impact of DitHub’s components, we progressively integrate them and present results in Figure˜7 (left).
 We establish the zero-shot performance of Grounding DINO (GD) as a reference.
@@
 Building on this baseline, we first add the warmup phase (+wu), which yields a notable improvement by stabilizing training and facilitating the learning of rare classes through robust initialization.
-Subsequently, incorporating Eq.
-3 into the warmup-augmented strategy results in a modest improvement — this is expected, as the shared $B$ matrix constitutes only a small fraction of the total parameters.
-Crucially, the most substantial performance gain is achieved upon integrating Eq.
-2 (DitHub in Fig.
-8), highlighting the importance of retaining domain-specific knowledge within class-specific modules to support stable and effective Incremental Learning.
+Subsequently, incorporating Equation˜4 into the warmup-augmented strategy results in a modest improvement — this is expected, as the shared $B$ matrix constitutes only a small fraction of the total parameters.
+Crucially, the most substantial performance gain is achieved upon integrating Equation˜3 (DitHub in Figure˜7 (left)), highlighting the importance of retaining domain-specific knowledge within class-specific modules to support stable and effective Incremental Learning.
 Collectively, the synergistic combination of these three key components – the warmup phase for robust initialization, the merging strategy for $A$ matrices to retain class-specific knowledge across domains, and the shared $B$ matrix for efficiency – yields a substantial $\boldsymbol{\texttt{+}6.13}$ mAP improvement compared to the baseline model.
+Finally, we note that results show a progressive decrease in the ZCOCO metric as the average mAP rises — an expected trade-off as specialization degrades zero-shot capabilities.
+Notably, the Base fine-tuning baseline scores higher on ZCOCO than the original Grounding DINO zero-shot model.
+We attribute this counter-intuitive result to catastrophic forgetting.
+Our full method, DitHub, is trained to be robust across the varied domains of ODinW.
+In contrast, the Base procedure completely forgets such broad knowledge and overfits to the new tasks.
+This process of forgetting diverse, out-of-distribution data results in a model that is coincidentally more aligned with the data distribution of the COCO benchmark, leading to an artificially inflated ZCOCO score.
+Analysis of the Performance-Memory Tradeoff.
+To study the trade-off between performance and memory, we examined how the average mAP on ODinW-13 is affected by the LoRA rank $r$, i.e., the dimensionality of the low-rank update matrices.
+The results, presented in Figure˜7 (right), demonstrate a graceful degradation in performance even as the rank is substantially reduced.
+Notably, a rank of $r=2$ achieves performance within $2$ mAP points of our best model ($r=16$) while reducing the per-class parameters and memory usage by a factor of eight.
+We also evaluated the extreme case of $r=1$, which collapses the adaptation matrix into a vector, and found that the performance remains remarkably strong.
+We attribute this robustness to our framework’s granularity; since each module only encodes the knowledge of a single class, it can operate effectively with fewer parameters than approaches that aggregate information from multiple classes into one module.
+In terms of memory efficiency, our method with $r=2$ requires approximately the same memory as ZiRa while performing significantly better ($\boldsymbol{\texttt{+}2.28}$ mAP).
+Furthermore, our $r=1$ variant delivers highly competitive results, scoring less than $1$ mAP point below ZiRa while using only half the memory.
+A detailed breakdown of the per-task mAP for these varying ranks is provided in Section˜A.3.
+Specialized modules allow training-free unlearning.
+The advantages of DitHub go beyond additive Incremental Learning.
+Expert modules can also be used to remove specific knowledge by simply subtracting the corresponding weights from the base model.
+This operation, akin to zero-training unlearning, has been studied in image and text categorization [15, 34]; to the best of our knowledge, we are the first to explore its application within the context of Object Detection.
+Such a mechanism allows the model to unlearn the detection of sensitive object categories – such as faces, license plates, or medical equipment – when required to comply with privacy regulations or data retention policies.
+Specifically, we test whether class knowledge can be unlearned from the zero-shot Grounding DINO baseline by subtracting the corresponding class $c$ module from the pre-trained model: $W^{\prime}=W-\alpha BA_{c}$, where we set $\alpha=0.3$.
+Figure˜8 qualitatively demonstrates DitHub’s class-unlearning capability.
+Each column illustrates the effect of removing a specific class module.
+For instance, when the horse module is subtracted, the previously detected horses are no longer recognized, resulting in false negatives that indicate successful forgetting.
+This highlights DitHub’s ability to selectively unlearn target classes while preserving knowledge of unrelated ones.
+A quantitative analysis is provided in Section˜D.2.
+Figure 8: Qualitative results of DitHub’s class unlearning. Each column compares zero-shot Grounding DINO predictions (left) with the output after subtracting a specific class module (right).
 6 Conclusions
@@
 DitHub will provide a fork of this manuscript to explore such questions.
+Acknowledgments and Disclosure of Funding
+Chiara Cappellino’s research was supported by the STORE project (https://edf-store.com), funded by the European Defence Fund (EDF) under grant agreement EDF-2022-101121405-STORE.
+Angelo Porrello was financially supported by the Italian Ministry for University and Research – through the ECOSISTER ECS 00000033 CUP E93C22001100001 project – and the European Commission under the Next Generation EU programme PNRR.
+Rita Cucchiara was financially supported by the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No.
+101120237).
 Appendix Overview
-This Appendix provides supplementary information regarding the main manuscript.
-  • App. A — Further Insights on Efficiency: Optimizes DitHub’s performance when adapting Grounding DINO.
-    • Sec. A.1 — Blockwise Adaptation Analysis: Justifies adapting only the encoder for DitHub based on empirical metrics.
-    • Sec. A.2 — Employing Full LoRA for Class-Specificity: Explores class-specific LoRA, favoring a parameter-efficient shared $B$ matrix.
-  • App. B — Further Insights on Forgetting: Analyzes DitHub’s anti-forgetting capabilities.
-    • Sec. B.1 — Forgetting on ODinW-13: Presents anti-forgetting results for ODinW-13.
-    • Sec. B.2 — Forgetting on ODinW-O: Demonstrates robust anti-forgetting on the ODinW-O benchmark.
-  • App. C — Additional Benchmark Details: Provides comprehensive dataset information.
-    • Sec. C.1 — ODinW-13 and MS COCO: Describes ODinW-13 for specialization and MS COCO (ZCOCO) for zero-shot evaluation.
-    • Sec. C.2 — ODinW-O: Introduces the ODinW-O benchmark for Incremental Learning and domain shifts.
-  • App. D — Additional Results: Presents complementary experimental information.
-    • Sec. D.1 — Standard Deviations: Reports standard deviations for reproduced methods from Tab. 1 and Tab. 2.
-    • Sec. D.2 — Unlearning Additionals: Illustrates DitHub’s modular capability for selective knowledge removal (Fig. C).
-  • App. E — Implementation Details: Outlines experimental setup, including hardware, model configurations, and optimization specifics.
-  • App. F — Limitations and Broader Impact: Discusses current limitations and broader societal implications.
-  • App. G — Qualitative Results: Provides visual comparisons (Fig. D) demonstrating DitHub’s improved detection and robustness.
+This appendix provides supplementary information to the main manuscript, organized as follows.
+Appendix A — Further Insights on Efficiency.
+Analyzes the performance optimization of DitHub when adapting Grounding DINO.
+  • Section˜A.1 — Blockwise Adaptation Analysis: Justifies adapting only the encoder for DitHub based on empirical metrics.
+  • Section˜A.2 — Employing Full LoRA for Class-Specificity: Explores class-specific LoRA, ultimately favoring a parameter-efficient shared $B$ matrix.
+  • Section˜A.3 — Performance-Memory Tradeo-ff Details: Details the per-task mAP for the experiments introduced in the right panel of Figure˜7.
+Appendix B — Further Insights on Forgetting.
+Evaluates the anti-forgetting capabilities of DitHub.
+  • Section˜B.1 — Forgetting on ODinW-13: Presents anti-forgetting results for the ODinW-13 benchmark.
+  • Section˜B.2 — Forgetting on ODinW-O: Demonstrates robust anti-forgetting performance on the ODinW-O benchmark.
+Appendix C — Additional Benchmark Details.
+Provides comprehensive information about the datasets used.
+  • Section˜C.1 — ODinW-13 and MS COCO: Describes the ODinW-13 dataset for specialization and MS COCO (ZCOCO) for zero-shot evaluation.
+  • Section˜C.2 — ODinW-O: Introduces the ODinW-O benchmark for Incremental Learning and domain shift analysis.
+Appendix D — Additional Results.
+Presents complementary experimental results.
+  • Section˜D.1 — Standard Deviations: Reports standard deviations for the reproduced methods in Table˜1 and Table˜2.
+  • Section˜D.2 — Further Unlearning Results: Illustrates the modular capability of DitHub for selective knowledge removal, as shown in Figure˜C.
+  • Section˜D.3 — Performance with Swin-B Backbone: Shows the scaling capabilities of DitHub when employing a larger backbone.
+  • Section˜D.4 — Beyond ODinW-13: Presents results for the AODRaw benchmark.
+Appendix E — Implementation Details.
+Outlines the experimental setup, including hardware, model configurations, and optimization specifics.
+Appendix F — Limitations and Broader Impact.
+Discusses the current limitations of the proposed method and its broader societal implications.
+Appendix G — Qualitative Results.
+Provides visual comparisons in Figure˜D that demonstrate the improved detection accuracy and robustness of DitHub.
 Table A: Performance comparison across datasets when adapting only the decoder (Only Dec), both encoder and decoder (Dec/Enc), and encoder only (DitHub).
@@
 This choice is motivated by the need for a more detection-centric adaptation and efficiency considerations, as the textual backbone, being a BERT Base model [5], accounts for more than half of Grounding DINO ’s parameters.
-In Tab.
-A, we present results for adapting only the encoder (Fig.
-A), only the decoder (Fig.
-A), and both encoder and decoder (Fig.
-A).
+In Table˜A, we present results for adapting only the encoder (Figure˜A), only the decoder (Figure˜A), and both encoder and decoder (Figure˜A).
 DitHub employs the first approach – adapting only the encoder – as it delivers the best performance on both ZCOCO and Avg metrics.
 A.2 Employing Full LoRA for Class-Specificity
-Table B: LoRA full results on ODinW-13, with DitHub results taken from Tab. 1 for reference.
+Table B: LoRA full results on ODinW-13, with DitHub results taken from Table˜1 for reference.
 Method ZCOCO Avg
@@
 DitHub (LoRA full) $46.66$ $63.22$
-While our methodology leverages a shared and unique $B$ matrix for all tasks, Tab.
-B presents results obtained using class-specific pairs.
+While our methodology leverages a shared and unique $B$ matrix for all tasks, Table˜B presents results obtained using class-specific pairs.
 In this variant of DitHub, both the $A$ and $B$ low-rank matrices of LoRA undergo the same specialized optimization process: an initial warmup phase followed by dedicated specialization.
-In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Sec.
-4.1.
-Applying our full set of strategies to both matrices yields a $\boldsymbol{\texttt{+}1.03}$ mAP improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Fig.
-3).
+In other words, the training pipeline for matrix $B$ mirrors that of matrix $A$ described in Section˜4.1.
+Applying our full set of strategies to both matrices yields a $\boldsymbol{\texttt{+}1.03}$ mAP improvement in average performance across the tasks of ODinW-13, albeit at the cost of doubling the number of trainable parameters (Figure˜3).
 Prioritizing efficiency, we opt for a lightweight implementation that trades a marginal performance gain for significantly reduced memory requirements.
 Notably, ZCOCO results favor our efficient approach, as a single, unspecialized $B$ matrix remains more aligned with the pretraining distribution, thereby mitigating degradation in zero-shot capabilities.
+A.3 Performance-Memory Trade-off Details
+Table C: Performance of DitHub at different ranks on the ODinW-13 benchmark.
+Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
+DitHub $r=1$ $47.14$ $57.04$ $31.89$ $37.01$ $67.93$ $72.58$ $43.63$ $53.76$ $69.80$ $67.86$ $42.92$ $72.49$ $40.31$ $76.11$ $65.27$
+DitHub $r=2$ $46.66$ $60.25$ $35.86$ $49.59$ $71.93$ $74.27$ $54.37$ $63.37$ $67.61$ $70.50$ $39.65$ $64.05$ $49.39$ $76.32$ $66.41$
+DitHub $r=4$ $46.73$ $60.53$ $38.54$ $48.59$ $71.13$ $74.45$ $50.46$ $61.49$ $68.20$ $69.97$ $41.48$ $64.97$ $55.64$ $78.13$ $63.84$
+DitHub $r=8$ $46.81$ $61.93$ $37.38$ $50.19$ $71.02$ $74.37$ $53.34$ $66.09$ $68.83$ $70.16$ $49.31$ $66.31$ $53.55$ $78.64$ $65.85$
+DitHub $r=16$ $47.01$ $62.19$ $34.62$ $50.65$ $70.46$ $68.56$ $49.28$ $65.57$ $69.58$ $71.10$ $56.65$ $70.88$ $52.82$ $79.30$ $68.18$
+Table˜C details the per-task mAP for the experiments introduced in Figure˜7 (right).
 Appendix B Further Insights on Forgetting
-This section presents an additional analysis of the anti-forgetting properties of DitHub, complementing the study in Sec.
-5.2 by offering an alternative perspective.
+This section presents an additional analysis of the anti-forgetting properties of DitHub, complementing the study in Section˜5.2 by offering an alternative perspective.
 We report the forgetting metric for both ODinW-13 and ODinW-O, computed for DitHub and its primary competitor, ZiRa [4].
@@
 B.1 Forgetting on ODinW-13
-Table C: Forgetting metrics in terms of mAP for the ODinW-13 dataset. The numbers above each task indicate the order in which they were processed. Bold indicates best results.
-$11$ $10$ $2$ $13$ $12$ $1$ $7$ $6$ $3$ $4$ $9$ $8$ $5$
+Table D: Forgetting metrics in terms of mAP for the ODinW-13 dataset. The numbers above each task indicate the order in which they were processed. Bold indicates best results.
+$11$ $10$ 22 $13$ $12$ 11 77 66 33 44 99 88 55
 Method Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
-ZiRa -$\texttt{-}6.66$$\texttt{-}4.31$$\texttt{-}4.82$$\boldsymbol{\texttt{-}1.80}$ $0.00$$\texttt{-}2.09$$\texttt{-}8.71$$\texttt{-}5.80$$\texttt{-}7.76$$\texttt{-}12.63$$\texttt{-}19.19$$\boldsymbol{\texttt{-}1.82}$$\texttt{-}9.85$$\texttt{-}7.77$
-DitHub -$\boldsymbol{\texttt{-}1.73}$$\boldsymbol{\texttt{-}1.56}$$\boldsymbol{\texttt{-}1.69}$$\texttt{-}2.34$ $0.00$$\boldsymbol{\texttt{-}1.61}$ $\boldsymbol{~{}~{}3.62}$$\boldsymbol{\texttt{-}1.91}$ $\boldsymbol{~{}~{}1.54}$$\boldsymbol{\texttt{-}6.94}$$\boldsymbol{\texttt{-}7.27}$$\texttt{-}2.11$ $\boldsymbol{~{}~{}0.10}$$\boldsymbol{\texttt{-}2.34}$
-Tab.
-C presents the forgetting in terms of mAP points for the $13$ tasks of ODinW-13, with the average across all tasks reported as Avg.
+ZiRa -$\texttt{-}6.66$ -$\texttt{-}4.31$ -$\texttt{-}4.82$ -$\boldsymbol{\texttt{-}1.80}$ $0.00$ -$\texttt{-}2.09$ -$\texttt{-}8.71$ -$\texttt{-}5.80$ -$\texttt{-}7.76$ -$\texttt{-}12.63$ -$\texttt{-}19.19$ -$\boldsymbol{\texttt{-}1.82}$ -$\texttt{-}9.85$ -$\texttt{-}7.77$
+DitHub -$\boldsymbol{\texttt{-}1.73}$ -$\boldsymbol{\texttt{-}1.56}$ -$\boldsymbol{\texttt{-}1.69}$ -$\texttt{-}2.34$ $0.00$ -$\boldsymbol{\texttt{-}1.61}$ $\boldsymbol{~~3.62}$ -$\boldsymbol{\texttt{-}1.91}$ $\boldsymbol{~~1.54}$ -$\boldsymbol{\texttt{-}6.94}$ -$\boldsymbol{\texttt{-}7.27}$ -$\texttt{-}2.11$ $\boldsymbol{~~0.10}$ -$\boldsymbol{\texttt{-}2.34}$
+Table˜D presents the forgetting in terms of mAP points for the $13$ tasks of ODinW-13, with the average across all tasks reported as Avg.
 DitHub demonstrates strong performance on both average and individual tasks, consistently outperforming its competitor by a notable margin.
@@
 B.2 Forgetting on ODinW-O
-Table D: Forgetting metrics in terms of mAP for the ODinW-O dataset. Bold indicates best results.
-$5$ $6$ $1$ $2$ $4$ $3$
+Table E: Forgetting metrics in terms of mAP for the ODinW-O dataset. Bold indicates best results.
+55 66 11 22 44 33
 Method Avg Ae Hw Pv Sd Th Ve
-ZiRa -$\texttt{-}3.84$$\boldsymbol{\texttt{-}0.08}$ $0.00$$\texttt{-}9.67$$\texttt{-}10.66$$\texttt{-}0.98$$\texttt{-}1.62$
-DitHub -$\boldsymbol{\texttt{-}2.68}$$\texttt{-}4.75$ $0.00$$\boldsymbol{\texttt{-}2.74}$$\boldsymbol{\texttt{-}9.51}$ $\boldsymbol{~{}~{}1.74}$$\boldsymbol{\texttt{-}0.79}$
-Tab.
-D reports the forgetting metric results for the ODinW-O benchmark.
+ZiRa -$\texttt{-}3.84$ -$\boldsymbol{\texttt{-}0.08}$ $0.00$ -$\texttt{-}9.67$ -$\texttt{-}10.66$ -$\texttt{-}0.98$ -$\texttt{-}1.62$
+DitHub -$\boldsymbol{\texttt{-}2.68}$ -$\texttt{-}4.75$ $0.00$ -$\boldsymbol{\texttt{-}2.74}$ -$\boldsymbol{\texttt{-}9.51}$ $\boldsymbol{~~1.74}$ -$\boldsymbol{\texttt{-}0.79}$
+Table˜E reports the forgetting metric results for the ODinW-O benchmark.
 Unlike the previous analysis, this assessment allows us to evaluate the forgetting of DitHub’s class-specific $A$ components for all classes within the dataset.
@@
 By enforcing strictly overlapping classes, ODinW-O provides a challenging benchmark for evaluating Incremental Learning.
-Additionally, to rigorously evaluate the model robustness to domain shifts, we designed ODinW-O by explicitly selecting datasets that feature markedly different contexts (Fig.
-B), enabling clear insights into how effectively the model adapts to variations in visual appearance, environmental conditions, and imaging domains.
-We select the following datasets, already mentioned in Sec.
-C.1, and extracted only the overlapping classes:
+Additionally, to rigorously evaluate the model robustness to domain shifts, we designed ODinW-O by explicitly selecting datasets that feature markedly different contexts (Figure˜B), enabling clear insights into how effectively the model adapts to variations in visual appearance, environmental conditions, and imaging domains.
+We select the following datasets, already mentioned in Section˜C.1, and extracted only the overlapping classes:
   • Aerial Maritime Drone: boat and car;
@@
 D.1 Standard Deviations
-We report the standard deviations of the reproduced methods from Tab. 1 and Tab. 2 in Tab. E and F, respectively, computed over three runs with different random seeds.
-D.2 Unlearning Additionals
-As previously discussed in Sec.
-5.2, our modular framework facilitates the precise removal of specific knowledge from Grounding DINO through the subtraction of corresponding experts.
-Fig.
-C visually represents this process: each column depicts the mean Average Precision (mAP) of a target class after unlearning different modules (reported on the y-axis), with colors normalized column-wise to emphasize relative performance decrements.
+We report the standard deviations of the reproduced methods from Table˜1 and Table˜2 in Table˜F and G, respectively, computed over three runs with different random seeds.
+D.2 Further Unlearning Results
+As previously discussed in Section˜5.2, our modular framework facilitates the precise removal of specific knowledge from Grounding DINO through the subtraction of corresponding experts.
+Figure˜C visually represents this process: each column depicts the mean Average Precision (mAP) of a target class after unlearning different modules (reported on the y-axis), with colors normalized column-wise to emphasize relative performance decrements.
 The prominent darker diagonal pattern signifies that removing a class-specific module results in the most substantial performance reduction for its associated class.
@@
 This correlation arises from the frequent co-occurrence of “person” instances riding bicycles, naturally linking these classes within the comprehensive unlearning pipeline.
-Table E: Standard deviations of the reproduced methods in Tab. 1.
+Table F: Standard deviations of the reproduced methods in Table˜1.
 Shots Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
-1 ZiRa ±$\pm 0.90$$\pm 2.05$$\pm 1.49$$\pm 0.17$$\pm 2.16$$\pm 0.98$$\pm 0.79$$\pm 1.63$$\pm 1.33$$\pm 7.91$$\pm 3.29$$\pm 12.17$$\pm 4.45$$\pm 3.07$$\pm 3.25$
-DitHub ±$\pm 0.60$$\pm 0.81$$\pm 0.95$$\pm 0.57$$\pm 2.66$$\pm 1.58$$\pm 3.72$$\pm 0.00$$\pm 1.53$$\pm 1.10$$\pm 2.90$$\pm 7.03$$\pm 3.68$$\pm 4.66$$\pm 2.46$
-5 ZiRa ±$\pm 0.23$$\pm 0.43$$\pm 3.76$$\pm 2.15$$\pm 1.40$$\pm 1.08$$\pm 3.54$$\pm 3.74$$\pm 1.54$$\pm 4.26$$\pm 3.73$$\pm 4.13$$\pm 2.03$$\pm 2.02$$\pm 4.16$
-DitHub ±$\pm 0.42$$\pm 1.27$$\pm 2.52$$\pm 0.70$$\pm 2.42$$\pm 6.74$$\pm 0.76$$\pm 3.27$$\pm 0.17$$\pm 5.49$$\pm 4.79$$\pm 2.22$$\pm 4.29$$\pm 1.81$$\pm 1.54$
-10 ZiRa ±$\pm 0.28$$\pm 0.66$$\pm 0.44$$\pm 2.64$$\pm 2.77$$\pm 0.52$$\pm 6.43$$\pm 0.97$$\pm 1.04$$\pm 2.17$$\pm 2.04$$\pm 3.90$$\pm 1.21$$\pm 3.71$$\pm 0.87$
-DitHub ±$\pm 0.20$$\pm 0.54$$\pm 2.26$$\pm 0.72$$\pm 4.23$$\pm 3.53$$\pm 1.35$$\pm 3.19$$\pm 0.30$$\pm 2.74$$\pm 0.34$$\pm 2.99$$\pm 1.01$$\pm 1.15$$\pm 1.27$
-Full ZiRA ±$\pm 0.44$$\pm 0.68$$\pm 2.73$$\pm 1.17$$\pm 1.39$$\pm 4.77$$\pm 7.26$$\pm 2.19$$\pm 1.33$$\pm 0.33$$\pm 4.14$$\pm 3.57$$\pm 2.44$$\pm 2.82$$\pm 3.13$
-DitHub ±$\pm 0.16$$\pm 0.44$$\pm 2.39$$\pm 1.00$$\pm 2.83$$\pm 0.48$$\pm 1.34$$\pm 0.60$$\pm 0.38$$\pm 1.68$$\pm 0.76$$\pm 1.57$$\pm 2.87$$\pm 0.69$$\pm 1.01$
-Table F: Standard deviations of the reproduced methods in Tab. 2.
+1 ZiRa $\pm 0.90$ $\pm 2.05$ $\pm 1.49$ $\pm 0.17$ $\pm 2.16$ $\pm 0.98$ $\pm 0.79$ $\pm 1.63$ $\pm 1.33$ $\pm 7.91$ $\pm 3.29$ $\pm 12.17$ $\pm 4.45$ $\pm 3.07$ $\pm 3.25$
+DitHub $\pm 0.60$ $\pm 0.81$ $\pm 0.95$ $\pm 0.57$ $\pm 2.66$ $\pm 1.58$ $\pm 3.72$ $\pm 0.00$ $\pm 1.53$ $\pm 1.10$ $\pm 2.90$ $\pm 7.03$ $\pm 3.68$ $\pm 4.66$ $\pm 2.46$
+5 ZiRa $\pm 0.23$ $\pm 0.43$ $\pm 3.76$ $\pm 2.15$ $\pm 1.40$ $\pm 1.08$ $\pm 3.54$ $\pm 3.74$ $\pm 1.54$ $\pm 4.26$ $\pm 3.73$ $\pm 4.13$ $\pm 2.03$ $\pm 2.02$ $\pm 4.16$
+DitHub $\pm 0.42$ $\pm 1.27$ $\pm 2.52$ $\pm 0.70$ $\pm 2.42$ $\pm 6.74$ $\pm 0.76$ $\pm 3.27$ $\pm 0.17$ $\pm 5.49$ $\pm 4.79$ $\pm 2.22$ $\pm 4.29$ $\pm 1.81$ $\pm 1.54$
+10 ZiRa $\pm 0.28$ $\pm 0.66$ $\pm 0.44$ $\pm 2.64$ $\pm 2.77$ $\pm 0.52$ $\pm 6.43$ $\pm 0.97$ $\pm 1.04$ $\pm 2.17$ $\pm 2.04$ $\pm 3.90$ $\pm 1.21$ $\pm 3.71$ $\pm 0.87$
+DitHub $\pm 0.20$ $\pm 0.54$ $\pm 2.26$ $\pm 0.72$ $\pm 4.23$ $\pm 3.53$ $\pm 1.35$ $\pm 3.19$ $\pm 0.30$ $\pm 2.74$ $\pm 0.34$ $\pm 2.99$ $\pm 1.01$ $\pm 1.15$ $\pm 1.27$
+Full ZiRA $\pm 0.44$ $\pm 0.68$ $\pm 2.73$ $\pm 1.17$ $\pm 1.39$ $\pm 4.77$ $\pm 7.26$ $\pm 2.19$ $\pm 1.33$ $\pm 0.33$ $\pm 4.14$ $\pm 3.57$ $\pm 2.44$ $\pm 2.82$ $\pm 3.13$
+DitHub $\pm 0.16$ $\pm 0.44$ $\pm 2.39$ $\pm 1.00$ $\pm 2.83$ $\pm 0.48$ $\pm 1.34$ $\pm 0.60$ $\pm 0.38$ $\pm 1.68$ $\pm 0.76$ $\pm 1.57$ $\pm 2.87$ $\pm 0.69$ $\pm 1.01$
+Table G: Standard deviations of the reproduced methods in Table˜2.
 Method ZCOCO Avg Ae Hw Pv Sd Th Ve
-ZiRa ±$\pm 0.25$$\pm 0.27$$\pm 1.76$$\pm 1.14$$\pm 1.35$$\pm 0.32$$\pm 1.75$$\pm 1.76$
-DitHub ±$\pm 0.24$$\pm 0.66$$\pm 2.39$$\pm 1.88$$\pm 0.76$$\pm 1.31$$\pm 0.46$$\pm 0.84$
+ZiRa $\pm 0.25$ $\pm 0.27$ $\pm 1.76$ $\pm 1.14$ $\pm 1.35$ $\pm 0.32$ $\pm 1.75$ $\pm 1.76$
+DitHub $\pm 0.24$ $\pm 0.66$ $\pm 2.39$ $\pm 1.88$ $\pm 0.76$ $\pm 1.31$ $\pm 0.46$ $\pm 0.84$
+D.3 Performance with Swin-B Backbone
+Table H: Performance Evaluation with the Swin-B backbone on ODinW-13.
+Method ZCOCO Avg Ae Aq Co Eg Mu Pa Pv Pi Po Ra Sh Th Ve
+GD ZS w/Swin-B $56.47$ $59.82$ $33.32$ $43.26$ $72.98$ $74.90$ $37.67$ $59.14$ $70.30$ $73.56$ $48.42$ $67.30$ $62.61$ $70.69$ $64.48$
+ZiRa w/Swin-B $53.22$ $65.09$ $40.71$ $54.16$ $73.32$ $73.97$ $47.25$ $67.03$ $72.39$ $73.29$ $59.83$ $73.05$ $61.98$ $82.59$ $66.63$
+DitHub w/Swin-B $53.82$ $67.03$ $42.41$ $53.74$ $74.53$ $75.18$ $53.28$ $66.05$ $74.04$ $74.22$ $59.25$ $76.51$ $67.63$ $82.85$ $71.67$
+To further assess the scalability of our approach, we conducted new experiments utilizing the larger, publicly available pre-trained Grounding DINO model with a Swin-B backbone.
+Table˜H details the comparative performance of our method, DitHub, against the primary competitor, ZiRa.
+The findings confirm that DitHub scales effectively, maintaining its robust performance when integrated with a more powerful backbone model.
+D.4 Beyond ODinW-13
+Table I: Domain-wise performance comparison on ODinW-13.
+Method ZCOCO Avg fog_low_light fog_rain fog indoor_low_light indoor_normal_light low_light_rain low_light normal_light rain
+GD ZS $47.40$ $15.65$ $13.32$ $19.56$ $19.42$ $15.84$ $24.77$ $11.74$ $16.28$ $5.81$ $14.11$
+ZiRA $44.39$ $32.34$ $27.60$ $36.34$ $29.28$ $28.28$ $43.18$ $30.67$ $31.73$ $30.49$ $33.47$
+DitHub $45.23$ $34.03$ $30.30$ $35.48$ $31.40$ $31.02$ $47.77$ $31.36$ $32.88$ $30.90$ $35.12$
+To rigorously evaluate the generalizability of our approach, we have extended our experiments to a more diverse benchmark.
+Our initial experimental setup necessitated a benchmark with class re-occurrence across semantically distinct domains — a characteristic not commonly found in standard long-tailed datasets.
+For this purpose, we selected AODRaw [22], a challenging and recently introduced benchmark.
+This dataset is ideally suited for our evaluation, as it comprises $9$ distinct domains with substantial class overlap, encompassing a wide range of meteorological conditions as well as both indoor and outdoor scenes.
+The performance analysis is presented in Table˜I.
 Appendix E Implementation Details
@@
 Appendix G Qualitative Results
-In Fig.
-D, we compare zero-shot Grounding DINO (left) with ZiRA (center) and DitHub (right).
+In Figure˜D, we compare zero-shot Grounding DINO (left) with ZiRA (center) and DitHub (right).
 Our method excels at detecting challenging objects, including small ones that other models miss.